### v4

# Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control

Abstract
The advanced bidirectional EV charging and discharging technology, aimed at supporting grid stability and emergency operations, has driven a growing interest in workplace applications. It not only reduces electricity expenses but also enhances the resilience in handling practical matters, such as peak power limitation, fluctuating energy prices, and unpredictable EV departures. Considering these factors systematically can benefit energy efficiency in office buildings and for EV users simultaneously. To employ AI to address these issues, we propose HUCA, a novel real-time charging control for regulating energy demands for both the building and EVs. HUCA employs hierarchical actor-critic networks to dynamically reduce electricity costs in buildings, accounting for the needs of EV charging in the dynamic pricing scenario. To tackle the uncertain EV departures, we introduce a new critic augmentation to account for departure uncertainties in evaluating the charging decisions, while maintaining the robustness of the charging control. Experiments on real-world electricity datasets under both simulated certain and uncertain departure scenarios demonstrate that HUCA outperforms baselines in terms of total electricity costs while maintaining competitive performance in fulfilling EV charging requirements. A case study also manifests that HUCA effectively balances energy supply between the building and EVs based on real-time information, showcasing its potential as a key AI-driven solution for vehicle charging control.
keywords: Hierarchical reinforcement learning, EV bidirectional charging, real-time charging control, dynamic pricing
1 Introduction
Background. The rapid growth in Electric Vehicle (EV) adoption, driven by global sustainability efforts, has led to a surge in demand for advanced charging solutions. According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [4], workplace charging sessions increased at twice the rate of new charging station installations in 2023. This highlights the growing interest for EV charging options in the workplace. Meanwhile, recent advances in bidirectional charging systems [19] [23], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and emergency situations. The effectiveness of bidirectional charging systems has inspired us to explore how better scheduling of G2V and V2G timings can balance the demand for EV charging and reduce overall electricity costs.
Figure 1: The illustration of the workplace charging scenario with the bidirectional EV chargers.
In particular, EVs often remain stationary for long periods at office locations, as expected. This prolonged parking time allows for more flexible charging strategies, enabling energy transfer between EVs and office building. An example of a workplace charging system is shown in Figure 1. Upon arrival, each EV is coordinated by the charging system, with users first providing information such as their charging requirements and anticipated departure time. With such information, the system decides whether to discharge energy from the EV to the building (using V2G) during high-price periods or to charge the EV from the grid to meet its demand before departure (using G2V). Either the V2G or the G2V option can be determined on-the-fly according to the optimal decision-making criteria.
Challenges. For real-time charging control of EVs in various scenarios, previous studies [20] [6] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging and discharging actions. However, most existing approaches fail to consider real-world dynamic factors, such as dynamic energy prices and the possibility that EV users may depart earlier than the expected time, which complicate determining optimal control strategies for each EV. Moreover, to avert transformer overloads that could destabilize the power grid [24], it is necessary to impose charging power limits, thereby further complicating the management of EV charging. These dynamics and limitations pose significant challenges in balancing the energy supply between the building and EVs while minimizing electricity costs. It is crucial to recognize that managing charging improperly could result in considerably higher electricity bills, as power companies will levy extra charges due to overconsumption of energy [18].
Proposed Method. To tackle these challenges, we propose HUCA (Hierarchical Multi-Agent Control with Uncertainty-Aware Critic Augmentation), a novel framework designed for real-time charging control. HUCA operates under dynamic pricing and penalty mechanisms, consistent with practical situations. A new hierarchical multi-agent framework is devised to balance the energy supply between the building and EVs in dynamic and uncertain environments. The HUCA framework consists of two levels of control: a high-level agent and multiple low-level agents. The high-level agent determines whether to charge or discharge EVs, which is a collective decision for all EVs. Based on the high-level decision, the low-level agents collaboratively and dynamically modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload.
A key innovation of HUCA lies in its capability to handle EVs that deviate from their expected departure times. These unpredictable departures introduce uncertainty, potentially disrupting low-level agents’ charging decisions. To address this, we propose an uncertainty-aware critic augmentation mechanism, which assesses the likelihood of departure deviation and adjusts the assessments of low-level agents’ charging decisions accordingly. This mechanism achieves two objectives: (i) accounting for uncertainties into the assessment of low-level agents’ decisions, and (ii) limiting the direct effect of uncertainties to the present time slot, ensuring the robustness of future decision assessments. Therefore, HUCA dynamically adapts to arbitrary EV behaviors while maintaining robust charging control. HUCA is designed to minimize the electricity costs of the building while striving to fulfill EV charging demands in a dynamic environment.
The main contributions in this paper are as follows:
  • Real-Time Charging Control: We propose HUCA, a novel hierarchical multi-agent structure that balances energy supply between the building and EVs, which optimizes charging controls on-the-fly.
  • Uncertainties and Limitations Handling: High-level and low-level agents determine optimal actions under dynamic pricing and power stability constraints. In addition, an uncertainty-aware augmentation is designed to adaptively adjust decisions to handle uncertain EV departures.
  • Practical Effectiveness: We evaluate HUCA through analysis of real-world datasets, incorporating simulated EV behaviors in scenarios with both certain and uncertain departure times. Results demonstrate that HUCA achieves the lowest electricity costs while maintaining competitive performance in fulfilling EV charging requirements, which shows its potential as an AI-driven solution for intelligent vehicle charging control.
2 Preliminaries
Figure 2: The overview of the HUCA framework. The high-level agent decides to charge or discharge EVs, while low-level agents control the power of each charging pile within power limitations.
This section delineates the essential symbols and their definitions. We also introduce the general formulations of Markov decision process (MDP) and deep deterministic policy gradient (DDPG) related to reinforcement learning techniques.
2.1 Key Symbols and Definitions.
Let $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$ denote a set of charging piles (abbreviated as piles henceforth) located at the office building’s charging station. $\mathcal{V}_{t}=\{v_{1},v_{2},...,v_{M}\}$ specifies a set of EVs docking at piles at time slot $t$, where $M\leq N$. The charging capacity of the charging station is denoted as $\mathcal{P}^{\text{max}}$ (kW), which represents the maximum allowable charging power at the station ($\mathcal{P}^{\text{max}}>0$). In contrast, the maximum allowable discharging power is denoted as $-\mathcal{P}^{\text{max}}$ (kW).
Definition 1. (State of Charge (SoC)): The state of charge (SoC) of an EV battery represents the percentage of energy stored in the battery. For each EV $v_{i}\in\mathcal{V}_{t}$, the SoC of $v_{i}$ at time slot $t$ is denoted as $SoC^{i}_{t}\in[0\%,100\%]$.
Definition 2. (EV Charging Information): In our scenario, when an EV arrives at the workplace charging system (Fig. 1) at time slot $t$ and connects to pile $c_{i}$, the EV user sends the charging information $\mathcal{I}^{i}$ to the charging system. The charging information is denoted as a tuple: $\mathcal{I}^{i}=(t^{i}_{\text{arr}},t^{i}_{\text{dep}},SoC^{i}_{\text{arr}},SoC^{i}_{\text{dep}},C^{i})$, where $t^{i}_{\text{arr}}$ and $t^{i}_{\text{dep}}$ represent the arrival time and the planned departure time, respectively. $SoC^{i}_{\text{arr}}$ and $SoC^{i}_{\text{dep}}$ denote the SoC of the EV battery upon arrival and the expected SoC to be reached by the charging system before the departure time, respectively. $C^{i}$ is the battery capacity of the EV. Note that in our scenario, the actual departure time of the EV user may randomly occur earlier than the planned time $t^{i}_{\text{dep}}$.
Definition 3. (Charging Power Limitation): To ensure power system stability, the charging power of each charging pile is limited by the maximum charging capacity $\mathcal{P}^{\text{max}}$ of the charging station. Given the number of EVs $|\mathcal{V}_{t}|$ docking at the charging station at time slot $t$, the maximum charging power $P^{\text{pile}}_{t}$ of each pile at time slot $t$ is estimated as $P^{\text{pile}}_{t}=\mathcal{P}^{\text{max}}/|\mathcal{V}_{t}|$. The maximum discharging power of each pile is $-P^{\text{pile}}_{t}$.
To consider the fulfillment of the expected SoC $SoC^{j}_{\text{dep}}$ of each EV $v_{j}$, we estimate the specific charging/discharging power boundary that pile $c_{i}$ provides to the connected EV $v_{j}$ at each time slot $t$, following the settings in  [9, 11]. Specifically, given the minimum and maximum SoC that EV $v_{j}$ can reach at time $t$, denoted by $SoC^{LB}_{t}$ and $SoC^{UB}_{t}$, respectively, the specific charging and discharging power boundaries of charging pile $c_{i}$ are as follows:
$\left\{\begin{matrix}\mathcal{P}^{\max}_{i,t}=\min\bigl{(}P^{\text{pile}}_{t},(SoC^{UB}_{t}-SoC^{j}_{t-1}\cdot C^{i}\cdot\eta^{*})\bigr{)}\\ \mathcal{P}^{\min}_{i,t}=\max\bigl{(}-P^{\text{pile}}_{t},(SoC^{LB}_{t}-SoC^{j}_{t-1}\cdot C^{j}\cdot\eta^{*})\bigr{)}\end{matrix}\right.,$
where $SoC^{j}_{t-1}$ is the SoC of $v_{j}$ at time slot $t-1$, and $C^{j}$ denotes the battery capacity for the EV. $\eta^{*}$ is function of charging efficiency, as defined in  [9, 11].
2.2 Markov Decision Process and Deep Deterministic Policy Gradient.
Decision-making problems are commonly modeled as Markov decision processes (MDPs), which defined by the four-tuple information $(S,A,R,T)$: states $S$, actions $A$, rewards $R$ and the state transition probabilities $T$. The main objective of an MDP is to find a policy that maximizes the cumulative rewards by selecting appropriate actions. The Deep Deterministic Policy Gradient (DDPG) algorithm [13], which is an actor-critic policy-based method that extends stochastic policy gradients to deterministic settings, has shown great effectiveness in complex environments. Specifically, its deep actor network, parameterized by $\theta^{\mu}$, approximates the deterministic policy $\mu_{\theta}:S\to A$. The critic network, parameterized by $\theta^{Q}$, estimates the action-value function $Q(s,a|\theta^{Q})$. Overall, the actor network determines the optimal action for a given state, while the critic network evaluates the action’s quality.
2.3 Problem Formulation.
Our work aims to benefit both the office building and EVs by enhancing the original single-agent DDPG framework. We model the problem as a hierarchical MDP to optimize multiple objectives simultaneously, aiming to find the most effective bidirectional charging strategy that reduces building overall electricity costs while considering EV charging needs.
At each time slot $t$, given the charging pile set $\mathcal{C}$, EV set $\mathcal{V}_{t}$, EV charging information $\{\mathcal{I}^{i}|v_{i}\in\mathcal{V}_{t}\}$, and the charging and discharging power boundaries $\{\mathcal{P}^{\max}_{i,t},\mathcal{P}^{\min}_{i,t}|v_{i}\in\mathcal{V}_{t}\}$, the objective is to determine the optimal charging (or discharging) power $\mathcal{P}^{\text{opt}}_{i,t}$ for each EV $v_{i}$ at time slot $t$, where $\mathcal{P}^{\min}_{i,t}\leq\mathcal{P}^{\text{opt}}_{i,t}\leq\mathcal{P}^{\max}_{i,t}$. By determining $\mathcal{P}^{\text{opt}}_{i,t}$, in the certain departure scenario, where the EV’s actual departure time matches the planned one, the goal is to minimize the total electricity cost of the building while ensuring that EV users’ expected SoCs are met at their departure. In scenarios where EVs depart earlier than expected [^1] When the departure occurs later than anticipated, achieving desired SoCs of these EVs is relatively straightforward compared to scenarios of early departure. Consequently, this paper concentrates on addressing the uncertainty of early departure., our objective is to achieve their desired SoCs to the greatest extent while minimizing costs, balancing the trade-off between cost reduction and maximizing EV charging fulfillment.
3 The HUCA Framework
The architecture of HUCA is illustrated in Figure 2. Initially, a high-level agent determines whether to charge or discharge EVs using real-time data (Sec. 3.1). Subsequently, multiple low-level agents manage the charging and discharging power for each individual pile, taking into account the unpredictability of EV departures (Sec. 3.2). Finally, HUCA can determine the optimal decision instantaneously (Sec. 3.3).
3.1 High-Level Charging Decision Making.
This section formulates a novel MDP for a high-level agent to decide whether to charge or discharge EVs.
3.1.1 State representations.
A state $s_{h}$, denoted by $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$, describes the status at the time slot $t$. The triplet includes: (i.) the electricity usage of the building and the electricity price, $I^{e}_{t}$; (ii.) EVs situations, $I^{v}_{t}$; and (iii.) the information learned by low-level agents at the previous time slot, $I^{l}_{t-1}$. Details are given in Appendix A.2.
3.1.2 Discrete actions for the high-level agent.
Let $a_{h}\in[0,1]$ denote the action selected by the high-level agent. It is then converted into the discrete high-level action $a^{\text{disc}}_{h}$ to determine whether to charge or discharge EVs, as defined below:
(3.1) $a^{\text{disc}}_{h}=\left\{\begin{matrix}1&\text{, if $a_{h}\geq 0.5$}\\ 0&\text{, if $a_{h}<0.5$}\end{matrix}\right.,$
where $a^{\text{disc}}_{h}=1$ represents charging, and $a^{\text{disc}}_{h}=0$ denotes discharging.
3.1.3 High-level objective and reward function.
The total electricity cost of a building is calculated based on two parts: (i) the total electricity consumption and (ii) the amount of electricity exceeding the contracted capacity [7], which is the upper bound of instant electricity consumption. The objective of the high-level agent is to minimize the total electricity cost.
Let $L_{t}$ represent the electricity load, $\Delta L_{t}$ denote the excess electricity consumption of the building, $p_{t}$ be the electricity price, and $t$ specify the time slot. The reward function $r_{h}(s_{h},a^{\text{disc}}_{h})$, briefed as $r_{h}$, is defined as:
(3.2) $r_{h}=\kappa\cdot\Bigl{(}\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-\sum_{t^{\prime}=1}^{t}\max(0,\Delta L_{t^{\prime}})\cdot\varphi}_{\text{electricity cost term}}\Bigr{)}$
$\hskip 11.38092pt+\underbrace{\Bigl{(}-|L_{t}-L_{\text{avg}}|\Bigl{)}}_{\text{electricity balance term}},$
where the first term represents the potential total electricity cost up to time $t$, weighted by an importance factor $\kappa\in[0,1]$. $\varphi\in\mathbb{R}$ is a fixed penalty coefficient weighting the exceeding instant electricity usage. The second term balances the electricity consumption compared to the previous average load $L_{\text{avg}}$. This design reduces the reward for costly or imbalanced charging, encouraging the high-level agent to avoid similar actions in future comparable states.
To evaluate charging or discharging actions, we apply DDPG concepts to update the high-level policy $\mu_{h}$. The high-level agent trains the critic network, with parameters $\theta^{Q}_{h}$, to approximate action-value function $Q_{h}(s_{h},a_{h})$ by minimizing the following loss:
(3.3) $\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl{[}\Bigl{(}Q_{h}(s_{h},a_{h})-y_{h}\Bigr{)}^{2}\biggr{]},$
$y_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\mu_{h}(s^{\prime}_{h})},$
where $\gamma$ is a discount factor. $\bar{Q}_{h}$ is the target Q-function used to stabilize learning. The replay buffer $\mathcal{D}_{h}$ stores the transition experiences of the high-level agent in the form of tuples $(s_{h},a_{h},r_{h},s^{\prime}_{h})$. Subsequently, the actor network $\theta^{\mu}_{h}$ updates the high-level policy $\mu_{h}$ via gradient descent as follows:
(3.4) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\!=\!\!\!\!\!\!\mathop{\mathbb{E}}_{s_{h},a_{h}\sim\mathcal{D}_{h}}\!\nabla_{\theta^{\mu}_{h}}\mu_{h}(a_{h}|s_{h})\nabla_{a_{h}}Q_{h}(s_{h},a_{h})|_{a_{h}=\mu_{h}(s_{h})}.$
3.2 Low-Level Control with Uncertainty-Aware Critic Augmentation.
Based on the charging or discharging decision made by the high-level action, the goal of the low-level control is to determine the optimal charging (or discharging) power level for each pile, taking into account the uncertainty of EV departures. Each pile $c_{i}\in\mathcal{C}$ is regarded as a low-level agent and its task is formulated as an MDP. A multi-agent structure including multiple piles is deployed for low-level control.
3.2.1 State representations.
For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ comprises two information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$: (i) the real-time charging information of the connected EV $v_{j}\in\mathcal{V}_{t}$, denoted as $I^{v_{j}}_{t}$; and (ii) the information learned by the high-level agents, denoted as $I^{h}_{t}$. For simplicity of the presentation, we provide additional details in Appendix A.2.
3.2.2 Low-level objective and the critic augmentation.
Multi-agent DDPG (MADDPG) [14] extends DDPG into a multi-agent policy gradient algorithm, where decentralized agents learn a centralized critic based on the observations and actions of all agents. Inspired by MADDPG, we formulate the objective of each critic network, $\theta^{Q,i}_{h}$, of a low-level agent $c_{i}\in\mathcal{C}$ by minimizing the following loss:
(3.5) $\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\biggl{[}\Bigl{(}Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})-y^{i}_{l}\Bigr{)}^{2}\biggr{]},$
$y^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},...,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\mu^{j}_{l}(s^{\prime j}_{l})},$
where $\textbf{x}=\{s^{1}_{l},s^{2}_{l},...,s^{N}_{l}\}$ collects the states of all low-level agents. $\mathcal{D}^{i}_{l}$ is the replay buffer of agent $c_{i}$, storing the transition experience with tuples $(s^{i}_{l},a^{i}_{l},r^{i}_{l},s^{{}^{\prime}i}_{l})$. $Q^{i}_{l}$ is the action-value function from the critic network of agent $c_{i}$, and $\bar{Q}^{i}_{l}$ is the target Q-function.
However, early departures of EVs occur occasionally, making the departure time uncertain and complicating the selection of the best low-level actions. To tackle this challenge, inspired by the Upper Confidence Bound (UCB) algorithm [1], which adjusts action-value estimations based on uncertainty (or confidence) and thereby affects the action decision, we propose a novel augmented action-value function, $\mathcal{Q}^{i}_{l}$, for the actor network of each low-level agent $c_{i}$. The idea is to discourage higher power level discharging actions when the given expected departure time is approaching and the expected SoC is still unmet. This augmentation incorporates uncertainty into the actor network for decision-making, while isolating direct impact of uncertainties from the critic network training, thereby maintaining the robustness of the critic network. Accordingly, $\mathcal{Q}^{i}_{l}$ is designed as follows:
(3.6) $\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})\cdot\biggl{(}1-$
$\hskip 51.21504pt\rho\cdot\underbrace{|\log_{2}(\bar{a}^{i}_{l}+\epsilon)|\cdot\sqrt{\frac{\max(\Delta SoC^{i}_{t},0)}{\Delta T^{i}_{t}}}}_{\text{the uncertainty factor}}\>\>\biggr{)}\Large\small\large,$
where $\bar{a}^{i}_{l}\in[0,1]$ is the action chosen by low-level agent $c_{i}$. A smaller $\bar{a}^{i}$ indicates a tendency for higher discharge power, as defined in Sec. 3.2. $\Delta SoC^{i}_{t}$ represents the difference between the expected SoC $SoC^{i}_{\text{dep}}$ and the current SoC $SoC^{i}_{t}$ of the EV connected to pile $c_{i}$. $\Delta T^{i}_{t}$ is the time difference between the current time $t$ and the planned departure time $t^{i}_{\text{dep}}$. $\rho\in\mathbb{R}$ is a fixed coefficient representing the impact of uncertainty, and $\epsilon$ is a small constant. Consequently, high power discharging action is penalized when the uncertainty factor increases.
Based on Eq. 3.6, the actor network $\theta^{\mu,i}_{l}$ of low-level agent $c_{i}$ updates the low-level policy $\mu^{i}_{l}$ with gradient descent as follows:
(3.7) $\nabla_{\theta^{\mu,i}_{l}}J(\mu^{i}_{l})=$
$\hskip 14.22636pt\mathop{\mathbb{E}}_{\textbf{x},\textbf{a}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\bigl{[}\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(a^{i}_{l}|s^{i}_{l})\nabla_{a^{i}_{h}}\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})|_{a^{i}_{l}=\mu^{i}_{l}(s^{i}_{l})}\bigr{]}.$
3.2.3 Continuous actions and the reward function for low-level agents.
Let $a^{i}_{l}\in[0,1]$ be the action selected by a low-level agent $c_{i}$. To entangle low-level action space with the high-level discrete decision $a^{\text{disc}}_{h}$, the continuous action $\bar{a}^{i}_{l}$ is formulated as follows:
(3.8) $\bar{a}^{i}_{l}=\left\{\begin{matrix}\delta+\delta\cdot\sigma(a^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=1$}\\ \delta\cdot\sigma(a^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=0$}\end{matrix}\right.,$
where $\sigma(\cdot)$ is the sigmoid function, and $\delta$ is set to 0.5. In this setting, the action space of $a^{i}_{l}$ is constrained to $[0,0.5)$ while the discrete high-level action is “discharging” ($a^{\text{disc}}_{h}=0$). In contrast, $a^{i}_{l}$ is limited to $[0.5,1]$ when the high-level action is “charging” ($a^{\text{disc}}_{h}=1$).
Subsequently, the optimal charging/discharging power $\mathcal{P}^{\text{opt}}_{i,t}$ provided by low-level agent (pile) $c_{i}\in\mathcal{C}$ to the connected EV at time slot $t$ is estimated based on the continuous low-level action $a^{i}_{l}$, as defined below:
(3.9) $\mathcal{P}^{\text{opt}}_{i,t}=\bar{a}^{i}_{l}\cdot(\mathcal{P}^{\max}_{i,t}-\mathcal{P}^{\min}_{i,t})+\mathcal{P}^{\min}_{i,t}.$
Based on the selected action $a^{i}_{t}$ and the optimal power decision $\mathcal{P}^{\text{opt}}_{i,t}$ at each time slot, the reward function $r_{l}(s^{i}_{l},a^{i}_{l})$ (abbreviated as $r^{i}_{l}$) for the low-level agent (pile) $c_{i}$ is formulated based on the charging cost at the current time slot $t$ and the difference between the current SoC and the expected SoC at the EV’s departure connected to $c_{i}$, defined as follows:
(3.10) $r_{l}=\omega\cdot(-\mathcal{P}^{\text{opt}}_{i,t}\cdot p_{t})+(-|SoC^{i}_{t}-SoC^{i}_{\text{dep}}|),$
where $p_{t}$ and $SoC^{i}_{t}$ are the energy price and the SoC of the EV at time slot $t$, respectively. $SoC^{i}_{\text{dep}}$ is the expected SoC at the EV’s departure. The parameter $\omega$ reflects the impact of the current charging cost.
3.3 Optimization for the Hierarchical Control.
Based on Sec. 3.1 and Sec. 3.2, the critic and actor networks are optimized with minibatches of samples from the replay buffers, which follows the soft update rule. Let $\theta_{h}=\{\theta^{Q}_{h},\theta^{\mu}_{h}\}$ and $\theta^{i}_{l}=\{\theta^{Q,i}_{l},\theta^{\mu,i}_{l}\}$ represent the parameters of target critic and actor networks of the high-level agent and the low-level agent ($c_{i}$), respectively. The parameters updates are as follows:
(3.11) $\theta^{\prime}_{h}\leftarrow\tau\theta_{h}+(1-\tau)\theta^{\prime}_{h},\quad\theta^{{}^{\prime}i}_{l}\leftarrow\tau\theta^{i}_{l}+(1-\tau)\theta^{{}^{\prime}i}_{l},$
where $\tau\in[0,1]$ is the coefficient controlling the update rate. Through these optimizations, the decision-making processes of both high-level and low-level agents are refined without requiring future information. The details of optimization formulation are given in Appendix A.3.
4 Experimental Results
4.1 Dataset and Experimental Setup.
4.1.1 Dataset Description and Preprocessing.
For our workplace charging system scenario, we use the following datasets: (i) Building electricity demands: the CU-BEMS dataset [17] provides office building electricity consumption. (ii) Pricing mechanism: the ComEd APIs [5] supply real-time pricing data. Also, the penalty mechanism is designed based on [18]. The data is aggregated into hourly intervals. The training period is from July 1, 2018, to August 31, 2018, with testing in the following month. Details are in Appendix A.1.
4.1.2 Simulation of EV Behaviors.
The charging information $\mathcal{I}^{i}=(t^{i}_{\text{arr}},t^{i}_{\text{dep}},SoC^{i}_{\text{arr}},SoC^{i}_{\text{dep}},C^{i})$ (Def. 2) of the EV is modeled using truncated normal distributions by following [10]. Table 1 shows the setting of random variables and $C^{i}$ is set to 60 kWh. Two departure scenarios are examined:
  1. 1.
    Certain departure scenario: The actual departure time $\widehat{t}^{i}_{dep}$ is the same as the expected time $t^{i}_{dep}$.
  2. 2.
    Uncertain departure scenario: The actual departure time is randomly sampled earlier than the expected time, $\widehat{t}^{i}_{dep}\in[1,t^{i}_{dep})$.
Table 1: Random variables for EV information.
Information Distribution Boundaries
Arrival time $\mathcal{N}(9,1^{2})$ $7\leq t^{i}_{\text{arr}}\leq 12$
Expected departure time $\mathcal{N}(19,1^{2})$ $16\leq t^{i}_{\text{dep}}\leq 23$
SoC upon arrival $\mathcal{N}(0.4,0.1^{2})$ $0.3\leq SoC^{i}_{\text{arr}}\leq 0.6$
Expected SoC at departure $\mathcal{N}(0.8,0.1^{2})$ $0.6\leq SoC^{i}_{\text{dep}}\leq 0.9$
Table 2: Performance comparison (HUCA performance) with different number of charging piles (CP). Except of the “OPT” method, the best and second best results for “Penalty Cost” ($\downarrow$) and “Total Cost” ($\downarrow$) are in bold and underlined, respectively. Cells in gray highlight the results that are greater than or equal to the median value for “SoC Maintenance” ($\uparrow$), “SoC Fulfillment” ($\uparrow$) and “User Satisfaction” ($\uparrow$).
Certain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
Penalty Cost (USD) 10 0.00 3682.85 2516.84 522.28 354.84 354.82
Total Cost (USD) 5805.28 9563.31 8426.36 6431.31 6220.33 6215.42
Penalty Cost (USD) 20 0.00 7818.52 5643.15 2271.32 1869.38 1643.47
Total Cost (USD) 5860.05 13929.11 11798.68 8413.61 7952.69 7728.57
Uncertain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
Penalty Cost (USD) 10 0.00 4671.74 1742.35 529.18 448.52 355.84
Total Cost (USD) 5691.96 10398.77 7479.04 6214.53 6087.48 5993.78
SoC Fulfillment (%) 53.77 54.05 46.04 40.34 27.10 40.77
SoC Maintenance (%) 42.52 -10.58 129.01 -120.61 11.87 45.40
User Satisfaction (%) 48.14 21.73 87.52 -40.13 19.48 43.08
Penalty Cost (USD) 20 0.00 5010.81 2002.12 966.51 931.01 354.90
Total Cost (USD) 5760.05 10823.07 7815.99 6721.82 6562.80 6013.52
SoC Fulfillment (%) 56.06 55.27 47.94 48.88 40.34 42.44
SoC Maintenance (%) 32.77 12.56 0.90 91.96 -1.16 124.41
User Satisfaction (%) 44.41 33.91 24.42 70.42 19.59 83.42
Table 3: Ablation study results for the uncertain departure scenario. (SoC fulfillment and maintenance scores are abbreviated as “Ful.” and “Main.”, respectively.)
CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
10 HUCA 355.84 5993.78 40.77 % 45.40%
w/o C.A. 2015.62 7761.55 62.49% 124.01%
w/o H. 2707.71 8483.96 63.10% 31.91%
w/o Either 723.46 6393.26 44.53% -12.53%
CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
20 HUCA 354.90 6013.52 42.44% 124.41%
w/o C.A. 9362.87 15309.23 69.73% 209.70%
w/o H. 3461.37 9261.72 58.61% 7.69%
w/o Either 618.46 6381.87 53.26% 43.02%
4.1.3 Baselines.
We compare our HUCA with following baselines. OPT: An oracle with full knowledge of future information serves the optimal solution of minimal total cost and minimal penalty cost. We include the single-agent (DDPG [13]) and multi-agent reinforcement learning models (IQL [22], VDN[21], and MADDPG [14]). These details are provided in Appendix A.1.4.
4.1.4 Evaluation Metrics.
  • Penalty Cost (USD) measures the cost of exceeding the contracted capacity [7], which multiplies the penalty rate with the peak amount of electricity (check Appendix A.1 for details).
  • Total Cost (USD) is the sum of the basic electricity cost (electricity load $L_{t}$ multiplied by dynamic pricing $p_{t}$) and the penalty cost over the testing data.
In the certain departure scenario, all EV charging demands are met since all baselines are subject to the charging power limitations (Sec. 2.1, Def. 3) and no unexpected EV departures occur. In the uncertain departure scenario, how well the charging demands are fulfilled when unexpected EV departures happen needs examinations. Given $\widehat{SoC}^{i}_{dep}$ and $SoC^{i}_{dep}$ denoting the actual and expected SoC of EV $v_{i}$ at the time of departure, and $SoC^{i}_{arr}$ is the SoC of $v_{i}$ upon arrival at the parking lot, the metrics are defined as follows:
  • SoC Fulfillment (%) evaluates how well the expected SoC of EV $v_{i}$ is fulfilled, defined as: $\widehat{SoC}^{i}_{\text{dep}}/SoC^{i}_{\text{dep}}$.
  • SoC Maintenance (%) checks the fulfillment status at the midpoint ($\widehat{SoC}^{i}_{\text{mid}}$) between the arrival and the actual departure time, verifying whether the SoC is consistently maintained under the charging control. A drop below the arrival SoC would result in extra expenses billed to users. The metric is defined as: $(\widehat{SoC}^{i}_{\text{mid}}-SoC^{i}_{\text{arr}})/(\widehat{SoC}^{i}_{\text{dep}}-SoC^{i}_{\text{arr}})$, where $\widehat{SoC}^{i}_{\text{mid}}$ denotes the SoC at the midpoint.
  • User Satisfaction (%) represents the average of SoC fulfillment and SoC maintenance, indicating EV users’ satisfaction with the charging performance.
The performance of these three metrics is reported as the average for all EVs per parking session.
Figure 3: Performance of HUCA with different $\rho$ values with 10 piles in the uncertain departure scenario.
4.2 Comparison Results.
As shown in Table 2, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total costs that do not rely on future information for charging control across both departure scenarios. In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs. Although the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, it still achieves a comparable results to all the other baselines. Note that combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
Among the multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost. However, the user satisfaction scores of MADDPG are much lower than those of HUCA across different numbers of charging piles. While VDN achieves higher SoC fulfillment score than HUCA with 20 charging piles, its maintenance scores are significantly lower. With 10 charging piles, VDN’s maintenance score even drops to a negative value (-120.61%), indicating that under MADDPG and VDN control, EVs’ SoCs are not maintained at arrival levels. Therefore, EVs primarily serve as power providers rather than being effectively charged, resulting in additional expenses for users.
4.3 Ablation Study and Analysis of the Uncertainty Impact.
We conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o Either). The results are shown in Table 3. Compared to HUCA, removing either the critic augmentation or the high-level agent increases SoC fulfillment and maintenance scores. However, this also leads to a significant rise in both penalty and total costs. On the other hand, removing both components results in a slight increase in costs but causes a significant drop in SoC maintenance scores. These findings indicate that the hierarchical structure and critic augmentation in HUCA complement each other, and their combined use is essential for achieving the lowest costs while effectively balancing energy supply between the office building and EVs.
Furthermore, to assess the impact of the uncertainty term in HUCA for handling uncertain departures, Figure 3 compares total cost and the SoC fulfillment under various uncertainty coefficient $\rho$ (Eq. 3.6). As $\rho$ increases, total cost decreases significantly. Meanwhile, setting $\rho$ to its lowest value ($10^{-2}$) improves SoC fulfillment but comes at the cost of significantly higher total costs. These findings suggest that $\rho$ should not be set too low to ensure a better trade-off between cost and user charging demand. Properly tuning $\rho$ can further improve this balance, enhancing the overall effectiveness of HUCA.
4.4 Case Study.
To examine how HUCA balances energy supply between the building and EVs, Figure 4 visualizes the control decisions over a day. Two significant EV charging behaviors appear at 7:00 and 16:00 (red rectangle). At 7:00, both electricity demand and prices for the building are minimal (Figure 4(a)), therefore HUCA focuses on charging the EVs to reach their target SoC.At 16:00, even though there is a high demand for building energy and elevated prices, which typically would lead to transferring energy from the EVs to the building to reduce costs, HUCA opts to charge the EVs because of their high charging urgency (Figure 4(b)) using Eq. (3.6). In contrast, at 8:00 (green rectangle), HUCA discharges EVs to serve the high-usage building and avoid high pricing time simultaneously. These results manifest that HUCA effectively and dynamically balances energy supply between the building and EVs based on real-time information.
Figure 4: Details of the charging control for a day from 7:00 to 17:00. (a) shows building electricity demands and dynamic pricing; (b) visualizes average charging urgency, with $\Delta SoC$ as the difference between current and expected SoC, and $\Delta T$ as the remaining time until planned departure; (c) presents final control decisions.
5 Related Work
5.1 Real-time Charging Control.
Guo et al. [8] combined day-ahead optimization with real-time control to address uncertainties in solar output and EV parking behaviors. Chandra Mouli et al. [16] introduced an energy management system with vehicle-to-grid (V2G) services to adjust charging control. Chopra et al. [3] conducted a cost-benefit analysis for EV charging infrastructure in parking lots. Zhao et al. [29] employed risk-based day-ahead scheduling for EV aggregators using information gap decision theory to minimize building costs. Zhang et al. [28] developed an EV load optimization model to reduce grid-level energy costs by considering peak and valley electricity prices. Chen et al. [2] devised strategies for battery energy storage systems to participate in peak load shifting, minimizing load fluctuations.
5.2 Reinforcement Learning for Charging Management.
Single-agent RL, multi-agent RL (MARL), and hierarchical MARL approaches were used to manage charging systems. Single-agent RL represented the control system [12, 15] under various considerations, such as ensuring EV battery requirements [26] and addressing charger shortage [9]. MARL modeled fine-grained cooperations [25] among EVs and buildings to tackle safety concerns [27] and power transfer overload [6]. Hierarchical MARL further separated system operators and EV users, and coordinated them to minimize demand charges and energy costs [20].
Although literature in Sec. 5.1 and  5.2 optimized charging control with various aspects, most studies do not address the issue of uncertain EV departures. While some works[15, 8, 27] considered uncertainties in EV behavior, they primarily focused on a single objective, such as maximizing system revenue or minimizing user charging costs. In contrast, our work tackles dynamic factors, including fluctuating electricity prices and uncertain EV departures, aiming to simultaneously minimize total electricity costs for the charging system while accounting for EV users’ charging demands.
6 Conclusion
In this paper, we propose HUCA, a novel hierarchical multi-agent framework designed for real-time charging control in the workplace charging scenario. To address practical dynamics and limitations, HUCA integrates a hierarchical control and uncertainty-aware critic augmentation. With this design, HUCA adapts to dynamic factors, optimizes charging or discharging power for each charging pile, and accounts for departure uncertainties in evaluating the quality of charging control. Experiments demonstrate that, in both certain and uncertain departure scenarios, HUCA outperforms existing methods by achieving the lowest electricity costs while maintaining competitive performance in meeting EV charging requirements. A case study further illustrates that HUCA effectively balances energy supply between the building and EVs using real-time information.
A Reproducibility
A.1 Detailed Experimental Settings.
A.1.1 Dataset Preprocessing.
We utilize the CU-BEMS dataset [17], which contains detailed energy consumption data from an office building, recorded at one-minute intervals from July 2018 to December 2019. For dynamic pricing (real-time pricing) data, we use the APIs [5] provided by ComEd (Commonwealth Edison), an electric utility company, offering real-time 5-minute electricity pricing data through its hourly pricing program. In our study, we aggregate the total energy consumption from all electrical appliances in the CU-BEMS dataset into hourly data to represent the building’s energy consumption in our workplace charging scenario. The dynamic pricing data obtained from ComEd APIs is similarly aggregated into hourly intervals. The training period covers July 1, 2018, to August 31, 2018, and the testing period runs from September 1, 2018, to September 30, 2018.
A.1.2 Penalty Mechanism.
For the design of the penalty mechanism, the penalty charge is typically two or three times the basic capacity rate [18]. We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption)[7, 18], with detailed settings shown in Table 4. Let the maximum instant electricity load during the testing period be $L_{max}$. If $L_{max}<C_{contract}$, then the penalty is set to zero. Given a threshold coefficient $10<\delta<$ 1, the penalty is calculated by dividing the excess electricity into two tiers:
  1. 1.
    Tier 1 (Up to 10% Overload): For overloads up to 10% of the contract capacity ($0<L_{max}\leq\delta\cdot C_{contract}$), the penalty charge is set to $2\cdot R_{base}$.
  2. 2.
    Tier 2 (Beyond 10% Overload): For overloads exceeding 10% of the contracted capacity ($L_{max}>\delta\cdot C_{contract}$), the penalty charge is set to $3\cdot R_{base}$ for the amount that surpasses the 10% threshold.
Table 4: Design of the penalty mechanism.
Parameter Value
Contract Capacity ($C_{contract}$) 700 kW
Basic Capacity Rate ($R_{base}$) \$15/kW/month
Exceed Electricity Amount Penalty Charge
Up to 10% over $C_{contract}$ $2\cdot R_{base}$
More than 10% over $C_{contract}$ $3\cdot R_{base}$
Based on this setting, the total penalty cost is estimated as follows:
(A.1) Penalty $\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
$\hskip 71.13188pt+3\cdot R_{base}\cdot\max(0,L_{max}-\delta\cdot C_{contract}).$
A.1.3 Hyperparamter Settings.
The charging capacity ($\mathcal{P}^{\max}$) is set to 150 (kW), the EV battery capacity $C^{i}$ (Def. 2) is 60 (kwh), and the charging efficiency is 95%. For training our HUCA model, we use 1,500 episodes, with a replay buffer capacity of 30,000 for each low-level agent, and a batch size of 1,024. The penalty coefficient $\varphi$ (Eq. 3.2) weight factors $\kappa$ (Eq. 3.2), $\omega$ (Eq. 3.10), and the importance of the uncertainty term $\rho$ (Eq. 3.6) are set to 0.1, 0.1, 0.5, and 10, respectively.
A.1.4 Baseline Settings.
To evaluate the performance of the proposed HUCA model, we compared it against several established baseline approaches:
  • OPT: Assumes complete knowledge of future information, including all EV charging schedules, building energy demands, and dynamic electricity prices. This baseline uses a linear programming optimization model to solve the charging control problem.
  • DDPG (Deep Deterministic Policy Gradient) [13]: An actor-critic algorithm designed for continuous control in single-agent environments.
  • IQL (Independent Q-Learning) [22]: A decentralized multi-agent reinforcement learning approach where each agent independently learns its own Q-function.
  • VDN (Value Decomposition Networks)[21]: A centralized value-based reinforcement learning method used in multi-agent systems.
  • MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [14]: Extends DDPG to multi-agent settings by incorporating centralized training and decentralized execution.
Note that for all models (except for OPT), the states, actions, and rewards are configured in the same manner as for our low-level agents, but without incorporating the information provided by the high-level agent network.
A.2 State representations of high-level and low-level MDPs.
A.2.1 High-level Agent.
A state $s_{h}$ in the MDP optimizied by the high-level agent includes three main information $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$ at current time slot $t$. The details are as follows:
  • $I^{e}_{t}$ includes the electricity load of the building and electricity price for the following periods: the current time slot, the average over the past $n$ hours, and the historical price at the same time and day of the week.
  • $I^{v}_{t}$ includes the current number of EVs at the charging station and the amount of energy provided to them.
  • $I^{l}_{t-1}$ includes the average critic value (from the action-value function) and the standard deviation of all low-level agents from the previous time slot.
A.2.2 Low-level Agent.
For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ can be described as comprising two main information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$. Details are:
  • $I^{v_{j}}_{t}$ includes the current SoC of EV $v_{j}$ and the suitable maximum and minimum charging powers ($\mathcal{P}^{\max}_{i,t}$ and $\mathcal{P}^{\min}_{i,t}$) of agent (charging pile) $c_{i}$ as defined in Eq. 2.1.
  • $I^{h}_{t}$ includes the discrete high-level action $a^{\text{disc}}_{h}$ determined at time slot $t$, and the critic value of state-action pair $(s_{h},a_{h})$ obtained from the high-level critic network $Q_{h}(s_{h},a_{h};\theta^{Q}_{h})$ at time slot $t$.
A.3 Optimization of HUCA.
We optimize the HUCA framwork using minibatches of samples from the replay buffers. For the high-level agent, given a minibatch of $\mathcal{M}$ samples from the replay buffer $\mathcal{D}_{h}$, the training loss of the critic network is rewritten as:
(A.2) $\mathcal{L}(\theta^{Q}_{h})=\frac{1}{\mathcal{M}}\sum_{m}\Bigl{(}Q_{h}(s^{m}_{h},a^{m}_{h})-y^{m}_{h}\Bigr{)}^{2}.$
The actor network is updated using the sampled policy gradient as:
(A.3) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\approx\frac{1}{\mathcal{M}}\sum_{m}\nabla_{\theta^{\mu}_{h}}\mu_{h}(a^{m}_{h}|s^{m}_{h})\nabla_{a_{h}}Q_{h}(s^{m}_{h},a^{m}_{h}).$
Similarly, the critic and actor networks of each low-level agent $c_{i}$ are also optimized using minibatches.
B Limitation and Future Work
Despite the strong performance of HUCA and its consideration of multiple dynamic factors, we have not yet accounted for user willingness to adjust their charging preferences to supply energy to the building. A promising direction for future work is to incorporate user willingness, along with the buy-and-sell energy behaviors between users and the system, into our operational framework. User willingness could be integrated into the low-level actor-critic networks of HUCA. In our current design, the departure uncertainty term is used to set a lower confidence bound for the action-value function of each low-level agent. When incorporating user willingness, an incentive mechanism could be developed to estimate the upper confidence bound of the action-value function, as user willingness and incentives might increase the effectiveness of charging control.
This would not only help identify which EV users have greater flexibility in participating in energy control but also allow the system to evaluate both the lower and upper bounds of the potential quality of low-level actions. This approach could improve the optimal charging and discharging power control, enhancing both user satisfaction and the overall effectiveness of the system.

### v5

@@
-Uncertainty-Aware Critic Augmentation for Hierarchical Multi-Agent EV Charging Control
+Towards Hierarchical Multi-Agent Decision-Making for Uncertainty-Aware EV Charging
 Abstract
-The advanced bidirectional EV charging and discharging technology, aimed at supporting grid stability and emergency operations, has driven a growing interest in workplace applications.
-It not only reduces electricity expenses but also enhances the resilience in handling practical matters, such as peak power limitation, fluctuating energy prices, and unpredictable EV departures.
-Considering these factors systematically can benefit energy efficiency in office buildings and for EV users simultaneously.
-To employ AI to address these issues, we propose HUCA, a novel real-time charging control for regulating energy demands for both the building and EVs.
-HUCA employs hierarchical actor-critic networks to dynamically reduce electricity costs in buildings, accounting for the needs of EV charging in the dynamic pricing scenario.
-To tackle the uncertain EV departures, we introduce a new critic augmentation to account for departure uncertainties in evaluating the charging decisions, while maintaining the robustness of the charging control.
-Experiments on real-world electricity datasets under both simulated certain and uncertain departure scenarios demonstrate that HUCA outperforms baselines in terms of total electricity costs while maintaining competitive performance in fulfilling EV charging requirements.
-A case study also manifests that HUCA effectively balances energy supply between the building and EVs based on real-time information, showcasing its potential as a key AI-driven solution for vehicle charging control.
-keywords: Hierarchical reinforcement learning, EV bidirectional charging, real-time charging control, dynamic pricing
+Recent advances in bidirectional EV charging and discharging systems have spurred interest in workplace applications.
+However, real-world deployments face various dynamic factors, such as fluctuating electricity prices and uncertain EV departure times, that hinder effective energy management.
+To address these issues and minimize building electricity costs while meeting EV charging requirements, we design a hierarchical multi-agent structure in which a high-level agent coordinates overall charge or discharge decisions based on real-time pricing, while multiple low-level agents manage individual power level accordingly.
+For uncertain EV departure times, we propose a novel uncertainty-aware critic augmentation mechanism for low-level agents that improves the evaluation of power-level decisions and ensures robust control under such uncertainty.
+Building upon these two key designs, we introduce HUCA, a real-time charging control framework that coordinates energy supply among the building and EVs.
+Experiments on real-world electricity datasets show that HUCA significantly reduces electricity costs and maintains competitive performance in meeting EV charging requirements under both simulated certain and uncertain departure scenarios.
+The results further highlight the importance of hierarchical control and the proposed critic augmentation under the uncertain departure scenario.
+A case study illustrates HUCA’s capability to allocate energy between the building and EVs in real time, underscoring its potential for practical use.
+keywords: Hierarchical reinforcement learning, Uncertain-aware control, EV bidirectional charging, Real-time charging
 1 Introduction
+Figure 1: The illustration of the workplace charging scenario with the bidirectional EV chargers.
 Background.
 The rapid growth in Electric Vehicle (EV) adoption, driven by global sustainability efforts, has led to a surge in demand for advanced charging solutions.
-According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [4], workplace charging sessions increased at twice the rate of new charging station installations in 2023.
+According to a report by CBRE Group (Coldwell Banker Richard Ellis Group) Incorporated [ev.demandFroChargingPile], workplace charging sessions increased at twice the rate of new charging station installations in 2023.
 This highlights the growing interest for EV charging options in the workplace.
-Meanwhile, recent advances in bidirectional charging systems [19] [23], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and emergency situations.
+Meanwhile, recent advances in bidirectional charging systems [sami2019bidirectional] [upputuri2023comprehensive], which support both Grid-to-Vehicle (G2V) and Vehicle-to-Grid (V2G) power flows, have shown the potential to enhance flexibility with regard to charging stability and response capabilities.
 The effectiveness of bidirectional charging systems has inspired us to explore how better scheduling of G2V and V2G timings can balance the demand for EV charging and reduce overall electricity costs.
-Figure 1: The illustration of the workplace charging scenario with the bidirectional EV chargers.
 In particular, EVs often remain stationary for long periods at office locations, as expected.
@@
 Challenges.
-For real-time charging control of EVs in various scenarios, previous studies [20] [6] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging and discharging actions.
+For real-time charging control of EVs in various scenarios, previous studies [saner2022cooperative] [Silva2020CoordinationOE] have explored the use of multi-agent reinforcement learning (MARL) techniques to regulate EV charging actions.
 However, most existing approaches fail to consider real-world dynamic factors, such as dynamic energy prices and the possibility that EV users may depart earlier than the expected time, which complicate determining optimal control strategies for each EV.
-Moreover, to avert transformer overloads that could destabilize the power grid [24], it is necessary to impose charging power limits, thereby further complicating the management of EV charging.
+Moreover, to avert transformer overloads [^1] Transformer overloads occur when the demand for electricity exceeds the capacity of a transformer, thereby increasing the risk of overheating and subsequent failure.
+that could destabilize the power grid [visakh2023analysis], it is necessary to impose charging power limits, thereby further complicating the management of EV charging.
 These dynamics and limitations pose significant challenges in balancing the energy supply between the building and EVs while minimizing electricity costs.
-It is crucial to recognize that managing charging improperly could result in considerably higher electricity bills, as power companies will levy extra charges due to overconsumption of energy [18].
+It is crucial to recognize that managing charging improperly could result in considerably higher electricity bills, as power companies will levy extra charges due to overconsumption of energy [rosado2020framework].
 Proposed Method.
@@
 A new hierarchical multi-agent framework is devised to balance the energy supply between the building and EVs in dynamic and uncertain environments.
-The HUCA framework consists of two levels of control: a high-level agent and multiple low-level agents.
+HUCA consists of two levels of control: a high-level agent and multiple low-level agents.
 The high-level agent determines whether to charge or discharge EVs, which is a collective decision for all EVs.
-Based on the high-level decision, the low-level agents collaboratively and dynamically modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload.
+Based on the high-level decision, the low-level agents collaboratively modulate the individual charging (or discharging) power level for each EV, ensuring they stay within the specified charging power limits, with the goal of satisfying anticipated demands and avoiding transformer overload.
 A key innovation of HUCA lies in its capability to handle EVs that deviate from their expected departure times.
@@
 This mechanism achieves two objectives: (i) accounting for uncertainties into the assessment of low-level agents’ decisions, and (ii) limiting the direct effect of uncertainties to the present time slot, ensuring the robustness of future decision assessments.
-Therefore, HUCA dynamically adapts to arbitrary EV behaviors while maintaining robust charging control.
+Therefore, HUCA dynamically adapts to arbitrary EV behaviors while maintaining robust control.
 HUCA is designed to minimize the electricity costs of the building while striving to fulfill EV charging demands in a dynamic environment.
 The main contributions in this paper are as follows:
-  • Real-Time Charging Control: We propose HUCA, a novel hierarchical multi-agent structure that balances energy supply between the building and EVs, which optimizes charging controls on-the-fly.
+  • Real-Time Charging Control: We propose HUCA, a novel hierarchical multi-agent structure that balances energy supply between the building and EVs, optimizing charging controls on-the-fly.
   • Uncertainties and Limitations Handling: High-level and low-level agents determine optimal actions under dynamic pricing and power stability constraints.
@@
 Results demonstrate that HUCA achieves the lowest electricity costs while maintaining competitive performance in fulfilling EV charging requirements, which shows its potential as an AI-driven solution for intelligent vehicle charging control.
+2 Related Work
+2.1 Real-time Charging Control.
+Optimizing charging control considers different pricing mechanisms, such as time-of-use pricing [ting2024online, Zhang2023ElectricVL], real-time pricing [bitencourt2017optimal, selim2021electric], and so on.
+Also, Day-ahead optimization combined with real-time control has been used to address uncertainties in solar output and EV parking behaviors [Guo2016TwoStageEO], electricity prices [Zhao2017RiskBasedDS] as well as energy demand and PV generation [Yan2019OptimizedOC].
+Integrated management systems, additional energy resources or battery energy storage systems have also been developed for the adaptive control [Chen2020PowerCS, Chopra2023CostBA, ChandraMouli2019IntegratedPC].
+2.2 Reinforcement Learning for Charging Management.
+Single-agent RL, multi-agent RL (MARL), and hierarchical MARL approaches were used to manage charging systems.
+Single-agent RL represented the control system [Li2020ConstrainedEC] [Mhaisen2020RealTimeSF] under specific considerations, such as ensuring EV battery requirements [Zhang2021CDDPGAD], addressing charger shortage [Li2023OptimalEC], maximizing profits for charging stations [wang2019reinforcement], and minimizing users’ charging costs [Yan2021DeepRL].
+MARL modeled fine-grained cooperations [Silva2020CoordinationOE, Li2024MultiAgentGR, Park2022MultiagentDR, Yan2022ACC] among EVs and buildings to tackle safety concerns, power transfer overload [Silva2020CoordinationOE], and users’ charging cost or operating cost [Li2024MultiAgentGR, Park2022MultiagentDR, Yan2022ACC].
+Hierarchical MARL further separated system operators and EV users, and coordinated them to minimize demand charges and energy costs [saner2022cooperative].
+Although these studies optimized charging control with various aspects, most studies do not address the issue of uncertain EV departures.
+While some works [Mhaisen2020RealTimeSF, Guo2016TwoStageEO, Park2022MultiagentDR, wang2019reinforcement] consider uncertainties in EV departure, they primarily focused on a single objective, such as maximizing system revenue or minimizing user charging costs.
+Furthermore, in many of these studies, uncertainty factors are only encoded as part of the state information or leave them implicit in the stochastic transition dynamics in their RL models, without explicitly modeling or quantifying it, which limits the model’s ability to adaptively adjust charging control under varying levels of departure uncertainty.
+In contrast, our work addresses dynamic factors, including fluctuating electricity prices, EV charging demands, and the energy requirements of the office building.
+More importantly, we explicitly incorporate and quantify EV departure uncertainty in our model, allowing it to directly shape the reinforcement learning decision-making process, while avoiding current uncertainty from misleading future charging decisions.
+This enables reliable charging control that simultaneously minimizes total electricity costs and accounts for EV users’ charging demands in a dynamic environment.
 3 Preliminaries
 Figure 2: The overview of the HUCA framework. The high-level agent decides to charge or discharge EVs, while low-level agents control the power of each charging pile within power limitations.
-This section delineates the essential symbols and their definitions.
-We also introduce the general formulations of Markov decision process (MDP) and deep deterministic policy gradient (DDPG) related to reinforcement learning techniques.
+We describe key symbols, definitions, and general formulations of the Markov decision process (MDP) and deep deterministic policy gradient (DDPG) in reinforcement learning.
 3.1 Key Symbols and Definitions.
@@
 The charging capacity of the charging station is denoted as $\mathcal{P}^{\text{max}}$ (kW), which represents the maximum allowable charging power at the station ($\mathcal{P}^{\text{max}}>0$).
-In contrast, the maximum allowable discharging power is denoted as $-\mathcal{P}^{\text{max}}$ (kW).
 Definition 1.
@@
 Definition 3.
-(Charging Power Limitation): To ensure power system stability, the charging power of each charging pile is limited by the maximum charging capacity $\mathcal{P}^{\text{max}}$ of the charging station.
+(Charging Power Limitation): To ensure power system stability, the charging power of each charging pile is limited by the maximum charging capacity $\mathcal{P}^{\text{max}}$ of the charging station, following the settings in  [Li2023OptimalEC] [li2023constrained].
 Given the number of EVs $|\mathcal{V}_{t}|$ docking at the charging station at time slot $t$, the maximum charging power $P^{\text{pile}}_{t}$ of each pile at time slot $t$ is estimated as $P^{\text{pile}}_{t}=\mathcal{P}^{\text{max}}/|\mathcal{V}_{t}|$.
-The maximum discharging power of each pile is $-P^{\text{pile}}_{t}$.
-To consider the fulfillment of the expected SoC $SoC^{j}_{\text{dep}}$ of each EV $v_{j}$, we estimate the specific charging/discharging power boundary that pile $c_{i}$ provides to the connected EV $v_{j}$ at each time slot $t$, following the settings in  [9, 11].
-Specifically, given the minimum and maximum SoC that EV $v_{j}$ can reach at time $t$, denoted by $SoC^{LB}_{t}$ and $SoC^{UB}_{t}$, respectively, the specific charging and discharging power boundaries of charging pile $c_{i}$ are as follows:
-$\left\{\begin{matrix}\mathcal{P}^{\max}_{i,t}=\min\bigl{(}P^{\text{pile}}_{t},(SoC^{UB}_{t}-SoC^{j}_{t-1}\cdot C^{i}\cdot\eta^{*})\bigr{)}\\ \mathcal{P}^{\min}_{i,t}=\max\bigl{(}-P^{\text{pile}}_{t},(SoC^{LB}_{t}-SoC^{j}_{t-1}\cdot C^{j}\cdot\eta^{*})\bigr{)}\end{matrix}\right.,$
-where $SoC^{j}_{t-1}$ is the SoC of $v_{j}$ at time slot $t-1$, and $C^{j}$ denotes the battery capacity for the EV. $\eta^{*}$ is function of charging efficiency, as defined in  [9, 11].
+The maximum discharging power of each pile is $P^{\text{pile}}_{t}$.
+To consider the fulfillment of the expected SoC $SoC^{j}_{\text{dep}}$ of each EV $v_{j}$, we estimate the specific charging/discharging power boundary that pile $c_{i}$ provides to the connected EV $v_{j}$ at each time slot $t$.
+Specifically, given the minimum and maximum SoC that EV $v_{j}$ can reach at time $t$, denoted by $SoC^{LB}_{t}$ and $SoC^{UB}_{t}$, respectively, the specific charging and discharging power boundaries of pile $c_{i}$ are formulated according to [Li2023OptimalEC][li2023constrained] as:
+[…]
+where $SoC^{j}_{t-1}$ is the SoC of $v_{j}$ at time t1t-1, and $C^{j}$ denotes the battery capacity for $v_{j}$. $\eta^{*}$ is function of charging efficiency.
 3.2 Markov Decision Process and Deep Deterministic Policy Gradient.
-Decision-making problems are commonly modeled as Markov decision processes (MDPs), which defined by the four-tuple information $(S,A,R,T)$: states $S$, actions $A$, rewards $R$ and the state transition probabilities $T$.
+Decision-making problems are commonly modeled as MDPs, defined by the four-tuple information $(S,A,R,T)$: states $S$, actions $A$, rewards $R$, and the state transition probabilities $T$.
 The main objective of an MDP is to find a policy that maximizes the cumulative rewards by selecting appropriate actions.
-The Deep Deterministic Policy Gradient (DDPG) algorithm [13], which is an actor-critic policy-based method that extends stochastic policy gradients to deterministic settings, has shown great effectiveness in complex environments.
+The Deep Deterministic Policy Gradient (DDPG) algorithm [lillicrap2015continuous], which is an actor-critic policy-based method that extends stochastic policy gradients to deterministic settings, has shown great effectiveness in complex environments.
 Specifically, its deep actor network, parameterized by $\theta^{\mu}$, approximates the deterministic policy $\mu_{\theta}:S\to A$.
@@
 We model the problem as a hierarchical MDP to optimize multiple objectives simultaneously, aiming to find the most effective bidirectional charging strategy that reduces building overall electricity costs while considering EV charging needs.
-At each time slot $t$, given the charging pile set $\mathcal{C}$, EV set $\mathcal{V}_{t}$, EV charging information $\{\mathcal{I}^{i}|v_{i}\in\mathcal{V}_{t}\}$, and the charging and discharging power boundaries $\{\mathcal{P}^{\max}_{i,t},\mathcal{P}^{\min}_{i,t}|v_{i}\in\mathcal{V}_{t}\}$, the objective is to determine the optimal charging (or discharging) power $\mathcal{P}^{\text{opt}}_{i,t}$ for each EV $v_{i}$ at time slot $t$, where $\mathcal{P}^{\min}_{i,t}\leq\mathcal{P}^{\text{opt}}_{i,t}\leq\mathcal{P}^{\max}_{i,t}$.
+At each time slot $t$, given the charging pile set $\mathcal{C}$, EV set $\mathcal{V}_{t}$, EV charging information $\{\mathcal{I}^{j}|v_{j}\in\mathcal{V}_{t}\}$, and the charging and discharging power boundaries $\{\mathcal{P}^{\max}_{i,t},\mathcal{P}^{\min}_{i,t}|c_{i}\in\mathcal{C}\}$, the objective is to determine the optimal charging (or discharging) power $\mathcal{P}^{\text{opt}}_{i,t}$ for the EV connected to pile $c_{i}$ at time slot $t$, where $\mathcal{P}^{\min}_{i,t}\leq\mathcal{P}^{\text{opt}}_{i,t}\leq\mathcal{P}^{\max}_{i,t}$.
 By determining $\mathcal{P}^{\text{opt}}_{i,t}$, in the certain departure scenario, where the EV’s actual departure time matches the planned one, the goal is to minimize the total electricity cost of the building while ensuring that EV users’ expected SoCs are met at their departure.
-In scenarios where EVs depart earlier than expected [^1] When the departure occurs later than anticipated, achieving desired SoCs of these EVs is relatively straightforward compared to scenarios of early departure.
+In uncertain departure scenario, where EVs depart earlier than expected [^2] When the departure occurs later than anticipated, achieving desired SoCs of these EVs is relatively straightforward compared to scenarios of early departure.
 Consequently, this paper concentrates on addressing the uncertainty of early departure., our objective is to achieve their desired SoCs to the greatest extent while minimizing costs, balancing the trade-off between cost reduction and maximizing EV charging fulfillment.
@@
 Initially, a high-level agent determines whether to charge or discharge EVs using real-time data (Sec.
-3.1).
+4.1).
 Subsequently, multiple low-level agents manage the charging and discharging power for each individual pile, taking into account the unpredictability of EV departures (Sec.
-3.2).
+4.2).
 Finally, HUCA can determine the optimal decision instantaneously (Sec.
-3.3).
+4.3).
 4.1 High-Level Charging Decision Making.
 This section formulates a novel MDP for a high-level agent to decide whether to charge or discharge EVs.
-3.1.1 State representations.
-A state $s_{h}$, denoted by $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$, describes the status at the time slot $t$.
-The triplet includes: (i.) the electricity usage of the building and the electricity price, $I^{e}_{t}$; (ii.) EVs situations, $I^{v}_{t}$; and (iii.) the information learned by low-level agents at the previous time slot, $I^{l}_{t-1}$.
-Details are given in Appendix A.2.
-3.1.2 Discrete actions for the high-level agent.
+State representations. A state $s_{h}$ represents the status at time slot $t$, consisting of three main information $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$:
+  • $I^{e}_{t}$ includes the electricity load of the building and electricity price for the following periods: the current time slot, the average over the past $n$ hours, and the historical price at the same time and day of the week.
+  • $I^{v}_{t}$ includes the current number of EVs at the charging station and the amount of energy provided to them.
+  • $I^{l}_{t-1}$ includes the average and the standard deviation of critic value of all low-level agents from at time t1t-1.
+Discrete actions for the high-level agent.
 Let $a_{h}\in[0,1]$ denote the action selected by the high-level agent.
@@
 where $a^{\text{disc}}_{h}=1$ represents charging, and $a^{\text{disc}}_{h}=0$ denotes discharging.
-3.1.3 High-level objective and reward function.
-The total electricity cost of a building is calculated based on two parts: (i) the total electricity consumption and (ii) the amount of electricity exceeding the contracted capacity [7], which is the upper bound of instant electricity consumption.
+High-level objective and reward function.
+The total electricity cost of a building is calculated based on two parts: (i) the total electricity consumption and (ii) the amount of electricity exceeding the contracted capacity [^3] Typically, the electricity provider establishes a maximum power limit for each subscriber to efficiently manage overall consumption.
+If the instant electricity consumption of a subscriber surpasses this agreed-upon capacity, they will incur additional penalties.
+[Fernndez2013CostOO].
 The objective of the high-level agent is to minimize the total electricity cost.
 Let $L_{t}$ represent the electricity load, $\Delta L_{t}$ denote the excess electricity consumption of the building, $p_{t}$ be the electricity price, and $t$ specify the time slot.
-The reward function $r_{h}(s_{h},a^{\text{disc}}_{h})$, briefed as $r_{h}$, is defined as:
-(3.2) $r_{h}=\kappa\cdot\Bigl{(}\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-\sum_{t^{\prime}=1}^{t}\max(0,\Delta L_{t^{\prime}})\cdot\varphi}_{\text{electricity cost term}}\Bigr{)}$
+Define $L_{t^{\prime}}]_{+}=\max(0,\Delta L_{t^{\prime}})$.
+The reward function, $r_{h}(s_{h},a^{\text{disc}}_{h})$, abbreviated as $r_{h}$, is formulated as:
+(4.2) $\hskip-2.84544ptr_{h}=\kappa\cdot\Bigl(\underbrace{-\sum_{t^{\prime}=1}^{t}p_{t^{\prime}}\cdot L_{t^{\prime}}-[L_{t^{\prime}}]_{+}\cdot\varphi}_{\text{electricity cost term}}\Bigr)$
 $\hskip 11.38092pt+\underbrace{\Bigl(-|L_{t}-L_{\text{avg}}|\Bigl)}_{\text{electricity balance term}},$
@@
 The high-level agent trains the critic network, with parameters $\theta^{Q}_{h}$, to approximate action-value function $Q_{h}(s_{h},a_{h})$ by minimizing the following loss:
-(3.3) $\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl{[}\Bigl{(}Q_{h}(s_{h},a_{h})-y_{h}\Bigr{)}^{2}\biggr{]},$
-$y_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\mu_{h}(s^{\prime}_{h})},$
+(4.3) $\hskip-2.84544pt\mathcal{L}(\theta^{Q}_{h})=\mathop{\mathbb{E}}_{s_{h},a_{h},r_{h},s^{\prime}_{h}\sim\mathcal{D}_{h}}\biggl[\Bigl(Q_{h}(s_{h},a_{h})-y_{h}\Bigr)^{2}\biggr],$
+$\hskip-2.84544pty_{h}=r_{h}+\gamma\bar{Q}_{h}(s^{\prime}_{h},a^{\prime}_{h})|_{a^{\prime}_{h}=\bar{\mu}_{h}(s^{\prime}_{h})},$
 where $\gamma$ is a discount factor.
-$\bar{Q}_{h}$ is the target Q-function used to stabilize learning.
+$s^{\prime}_{h}$ is the next state after taking action $a_{h}$.
+$\bar{Q}_{h}$ and $\bar{\mu}_{h}$ are the target action-value function and target policy, respectively, used to stabilize training.
 The replay buffer $\mathcal{D}_{h}$ stores the transition experiences of the high-level agent in the form of tuples $(s_{h},a_{h},r_{h},s^{\prime}_{h})$.
-Subsequently, the actor network $\theta^{\mu}_{h}$ updates the high-level policy $\mu_{h}$ via gradient descent as follows:
-(3.4) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\!=\!\!\!\!\!\!\mathop{\mathbb{E}}_{s_{h},a_{h}\sim\mathcal{D}_{h}}\!\nabla_{\theta^{\mu}_{h}}\mu_{h}(a_{h}|s_{h})\nabla_{a_{h}}Q_{h}(s_{h},a_{h})|_{a_{h}=\mu_{h}(s_{h})}.$
+Subsequently, the actor network, with parameters $\theta^{\mu}_{h}$, refines the high-level policy $\mu_{h}$ via gradient descent as follows:
+(4.4) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\!=\!\!\!\!\!\!\mathop{\mathbb{E}}_{s_{h}\sim\mathcal{D}_{h}}\!\nabla_{\theta^{\mu}_{h}}\mu_{h}(a_{h}|s_{h})\nabla_{a_{h}}Q_{h}(s_{h},a_{h})|_{a_{h}=\mu_{h}(s_{h})}.$
+Using mini-batch updates, the high-level agent’s training loss is computed over transition experiences drawn from the replay buffer $\mathcal{D}_{h}$.
+At each training step, given a mini-batch $\mathcal{B}_{h}$ of transitions $\bigl\{(s_{h}^{(b)},a_{h}^{(b)},r_{h}^{(b)},s_{h}^{(b)})\bigr\}_{b=1}^{|\mathcal{B}_{h}|}$ drawn from $\mathcal{D}_{h}$, the critic and actor networks, parameterized by $\theta^{Q}_{h}$ and $\theta^{\mu}_{h}$, are updated as shown in Algorithm 1.
+Algorithm 1 High-Level Agent Update
+1:procedure UpdateHigh($\mathcal{B}_{h}$, $\theta^{Q}_{h}$, $\theta^{\mu}_{h}$)
+2:  Update the critic network by minimizing: $\mathcal{L}(\theta^{Q}_{h})=\frac{1}{\mathcal{B}_{h}}\sum_{b}\left(Q_{h}(s^{(b)}_{h},a^{(b)}_{h})-y{(b)}_{h}\right)^{2}$
+3:  Update the actor netwok using: $\nabla_{\theta^{\mu}_{h}}J\approx\frac{1}{\mathcal{B}_{h}}\sum_{b}\nabla_{\theta^{\mu}_{h}}\mu_{h}(s^{(b)}_{h})\nabla_{a_{h}}Q_{h}(s^{(b)}_{h},a^{(b)}_{h})$
+4:end procedure
 4.2 Low-Level Control with Uncertainty-Aware Critic Augmentation.
 Based on the charging or discharging decision made by the high-level action, the goal of the low-level control is to determine the optimal charging (or discharging) power level for each pile, taking into account the uncertainty of EV departures.
-Each pile $c_{i}\in\mathcal{C}$ is regarded as a low-level agent and its task is formulated as an MDP.
+Each pile $c_{i}\in\mathcal{C}$ is considered as a low-level agent and its task is formulated as an MDP.
 A multi-agent structure including multiple piles is deployed for low-level control.
-3.2.1 State representations.
-For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ comprises two information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$: (i) the real-time charging information of the connected EV $v_{j}\in\mathcal{V}_{t}$, denoted as $I^{v_{j}}_{t}$; and (ii) the information learned by the high-level agents, denoted as $I^{h}_{t}$.
-For simplicity of the presentation, we provide additional details in Appendix A.2.
+State representations.
+For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ can be described as comprising two main information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$.
+Details are:
+  • $I^{v_{j}}_{t}$ includes the current SoC of EV $v_{j}$ and the suitable maximum and minimum charging powers ($\mathcal{P}^{\max}_{i,t}$ and $\mathcal{P}^{\min}_{i,t}$) of agent (charging pile) $c_{i}$ as defined in Eq.
+3.1.
+  • $I^{h}_{t}$ includes the discrete high-level action $a^{\text{disc}}_{h}$ determined at time slot $t$, and the critic value of state-action pair $(s_{h},a_{h})$ obtained from the high-level critic network $Q_{h}(s_{h},a_{h};\theta^{Q}_{h})$ at time slot $t$.
 4.2.1 Low-level objective and the critic augmentation.
-Multi-agent DDPG (MADDPG) [14] extends DDPG into a multi-agent policy gradient algorithm, where decentralized agents learn a centralized critic based on the observations and actions of all agents.
+Multi-agent DDPG (MADDPG) [lowe2017multi] extends DDPG into a multi-agent policy gradient algorithm, where decentralized agents learn a centralized critic based on the observations and actions of all agents.
 Inspired by MADDPG, we formulate the objective of each critic network, $\theta^{Q,i}_{h}$, of a low-level agent $c_{i}\in\mathcal{C}$ by minimizing the following loss:
-(3.5) $\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\biggl{[}\Bigl{(}Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})-y^{i}_{l}\Bigr{)}^{2}\biggr{]},$
-$y^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},...,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\mu^{j}_{l}(s^{\prime j}_{l})},$
-where $\textbf{x}=\{s^{1}_{l},s^{2}_{l},...,s^{N}_{l}\}$ collects the states of all low-level agents.
-$\mathcal{D}^{i}_{l}$ is the replay buffer of agent $c_{i}$, storing the transition experience with tuples $(s^{i}_{l},a^{i}_{l},r^{i}_{l},s^{{}^{\prime}i}_{l})$.
-$Q^{i}_{l}$ is the action-value function from the critic network of agent $c_{i}$, and $\bar{Q}^{i}_{l}$ is the target Q-function.
+(4.5) $\hskip-2.84544pt\mathcal{L}(\theta^{Q,i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a},\textbf{r},\textbf{x}^{\prime}\sim\mathcal{D}_{l}}\biggl[\Bigl(Q^{i}_{l}(\textbf{x},a^{1}_{l},.,a^{N}_{l})-y^{i}_{l}\Bigr)^{2}\biggr],$
+$\hskip-2.84544pty^{i}_{l}=r^{i}_{l}+\gamma\bar{Q}^{i}_{l}(\textbf{x}^{\prime},a^{\prime 1}_{l},.,a^{\prime N}_{l})|_{a^{\prime j}_{l}=\bar{\mu}^{j}_{l}(s^{\prime j}_{l})},$
+where $\textbf{x}=\{s^{1}_{l},s^{2}_{l},...,s^{N}_{l}\}$ collects the states of all low-level agents, and $\textbf{x}^{\prime}$ contains all the next states.
+$\mathcal{D}_{l}$ is the replay buffer shared between low-level agents, storing the transition experience with tuples $(\textbf{x},a^{1}_{l},...,a^{N}_{l},r^{1}_{l},...,r^{N}_{l},\textbf{x}^{\prime})$.
+$Q^{i}_{l}$ is the action-value function from the critic network of agent $c_{i}$.
+$\bar{Q}^{i}_{l}$ and $\bar{\mu}^{j}_{l}$ are the target action-value function and the target policy, respectively.
+Figure 3: The illustration of uncertainty-aware critic augmentation.
+For each low-level agent $c_{i}$, the action-value function is augmented based on the uncertainty factor (Eq.
+4.6).
+This augmentation guides the updated policy to make decisions with explicit consideration of EV departure uncertainty.
 However, early departures of EVs occur occasionally, making the departure time uncertain and complicating the selection of the best low-level actions.
-To tackle this challenge, inspired by the Upper Confidence Bound (UCB) algorithm [1], which adjusts action-value estimations based on uncertainty (or confidence) and thereby affects the action decision, we propose a novel augmented action-value function, $\mathcal{Q}^{i}_{l}$, for the actor network of each low-level agent $c_{i}$.
-The idea is to discourage higher power level discharging actions when the given expected departure time is approaching and the expected SoC is still unmet.
-This augmentation incorporates uncertainty into the actor network for decision-making, while isolating direct impact of uncertainties from the critic network training, thereby maintaining the robustness of the critic network.
-Accordingly, $\mathcal{Q}^{i}_{l}$ is designed as follows:
-(3.6) $\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},...,a^{N}_{l})\cdot\biggl{(}1-$
-$\hskip 51.21504pt\rho\cdot\underbrace{|\log_{2}(\bar{a}^{i}_{l}+\epsilon)|\cdot\sqrt{\frac{\max(\Delta SoC^{i}_{t},0)}{\Delta T^{i}_{t}}}}_{\text{the uncertainty factor}}\>\>\biggr{)}\Large\small\large,$
-where $\bar{a}^{i}_{l}\in[0,1]$ is the action chosen by low-level agent $c_{i}$.
-A smaller $\bar{a}^{i}$ indicates a tendency for higher discharge power, as defined in Sec.
-3.2.
-$\Delta SoC^{i}_{t}$ represents the difference between the expected SoC $SoC^{i}_{\text{dep}}$ and the current SoC $SoC^{i}_{t}$ of the EV connected to pile $c_{i}$.
+To tackle this challenge, inspired by the Upper Confidence Bound (UCB) algorithm [auer2002finite], which adjusts action-value estimations based on uncertainty (or confidence) and thereby affects the action decision, we propose a novel augmented action-value function, $\mathcal{Q}^{i}_{l}$, for the actor network of each low-level agent $c_{i}$.
+[…]
+Our uncertainty-aware critic augmentation is illustrated in Figure 3, and the augmented action-value function $\mathcal{Q}^{i}_{l}$ is designed as follows:
+(4.6) $\hskip-8.5359pt\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})=Q^{i}_{l}(\textbf{x},a^{1}_{l},.,a^{N}_{l})\cdot\biggl(1$-
+$\hskip 51.21504pt\rho\cdot\underbrace{|\log_{2}(a^{i}_{l}+\epsilon)|\cdot\sqrt{\frac{\max(\Delta SoC^{i}_{t},0)}{\Delta T^{i}_{t}}}}_{\text{the uncertainty factor}}\>\>\biggr)\Large\small\large,$
+where $a^{i}_{l}\in[0,1]$ is the action chosen by low-level agent $c_{i}$.
+A smaller $a^{i}$ indicates a tendency for higher discharge power.
+$\Delta SoC^{i}_{t}$ represents the difference between the expected SoC $SoC^{i}_{\text{dep}}$ and the current SoC $SoC^{i}_{t}$ of the EV connected to low-level agent (pile) $c_{i}$.
 $\Delta T^{i}_{t}$ is the time difference between the current time $t$ and the planned departure time $t^{i}_{\text{dep}}$.
@@
 Consequently, high power discharging action is penalized when the uncertainty factor increases.
-Based on Eq. 3.6, the actor network $\theta^{\mu,i}_{l}$ of low-level agent $c_{i}$ updates the low-level policy $\mu^{i}_{l}$ with gradient descent as follows:
-(3.7) $\nabla_{\theta^{\mu,i}_{l}}J(\mu^{i}_{l})=$
-$\hskip 14.22636pt\mathop{\mathbb{E}}_{\textbf{x},\textbf{a}\sim\bigcup_{c_{i}\in\mathcal{C}}\{\mathcal{D}^{i}_{l}\}}\bigl{[}\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(a^{i}_{l}|s^{i}_{l})\nabla_{a^{i}_{h}}\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})|_{a^{i}_{l}=\mu^{i}_{l}(s^{i}_{l})}\bigr{]}.$
-3.2.3 Continuous actions and the reward function for low-level agents.
-Let $a^{i}_{l}\in[0,1]$ be the action selected by a low-level agent $c_{i}$.
-To entangle low-level action space with the high-level discrete decision $a^{\text{disc}}_{h}$, the continuous action $\bar{a}^{i}_{l}$ is formulated as follows:
-(3.8) $\bar{a}^{i}_{l}=\left\{\begin{matrix}\delta+\delta\cdot\sigma(a^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=1$}\\ \delta\cdot\sigma(a^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=0$}\end{matrix}\right.,$
+Based on Eq. 4.6, the actor network of low-level agent $c_{i}$, with parameters $\theta^{\mu,i}_{l}$, updates the low-level policy $\mu^{i}_{l}$ with gradient descent as:
+(4.7) $\small\nabla_{\theta^{\mu,i}_{l}}J(\mu^{i}_{l})=\mathop{\mathbb{E}}_{\textbf{x},\textbf{a}\sim\mathcal{D}_{l}}\bigl[\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(a^{i}_{l}|s^{i}_{l})\nabla_{a^{i}_{h}}\mathcal{Q}^{i}_{l}(\textbf{x},\textbf{a})|_{a^{i}_{l}=\mu^{i}_{l}(s^{i}_{l})}\bigr].$}
+Similar to the high-level agent, each low-level agent is trained with mini-batch updates.
+At every training step, a low-level agent $c_{i}$ draws a mini-batch $\mathcal{B}_{l}$ of transitions $\bigl\{(\textbf{x}^{(b)},a^{1,(b)}_{l},...,a^{N,(b)}_{l},r^{1,(b)}_{l},...,r^{N,(b)}_{l},\textbf{x}^{\prime(b)})\bigr\}_{b=1}^{\mathcal{B}_{l}}$ from the low-level replay buffer $\mathcal{D}_{l}$.
+The corresponding critic and actor networks, parameterized by $\theta^{Q,i}_{l}$ and $\theta^{\mu,i}_{l}$, are then updated as described in Algorithm 2.
+Algorithm 2 Low-Level Agent Update
+1:procedure UpdateLow($\mathcal{B}_{l}$, $\theta^{Q,i}_{l}$, $\theta^{\mu,i}_{l}$)
+2:  Update the critic network by minimizing: $\mathcal{L}(\theta^{Q,i}_{l})=\frac{1}{\mathcal{B}_{l}}\sum_{b}\Big(Q^{i}_{l}(\textbf{x}^{(b)},a^{1,(b)}_{l},...,a^{N,(b)}_{l})-y^{i,(b)}_{l}\Big)^{2}$
+3:
+4:  Compute the augmented action-value function $\mathcal{Q}^{i}_{l}$ $\triangleright$ Eq. 4.6
+5:  Update the actor network based on $\mathcal{Q}^{i}_{l}$: $\nabla_{\theta^{\mu,i}_{l}}J\approx\frac{1}{\mathcal{B}_{l}}\sum_{b}\nabla_{\theta^{\mu,i}_{l}}\mu^{i}_{l}(s^{i,(b)}_{l})\nabla_{a^{i,(b)}_{l}}\mathcal{Q}^{i}_{l}(\textbf{x}^{(b)},\textbf{a}^{(b)})$
+6:end procedure
+Continuous actions and rewards for low-level agents.
+In Eq.
+4.6, the action $a^{i}_{l}$ is derived from the raw logits $g^{i}_{l}$ produced by the low-level policy $\mu^{i}_{l}$ of agent $c_{i}$.
+To further entangle low-level action space with the high-level discrete decision $a^{\text{disc}}_{h}$, the continuous action $a^{i}_{l}$ is formulated as follows:
+(4.8) $a^{i}_{l}=\left\{\begin{matrix}\delta+\delta\cdot\sigma(g^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=1$}\\ \delta\cdot\sigma(g^{i}_{l})&\text{, if $a^{\text{disc}}_{h}=0$}\end{matrix}\right$.,
 where $\sigma(\cdot)$ is the sigmoid function, and $\delta$ is set to 0.5.
@@
 Subsequently, the optimal charging/discharging power $\mathcal{P}^{\text{opt}}_{i,t}$ provided by low-level agent (pile) $c_{i}\in\mathcal{C}$ to the connected EV at time slot $t$ is estimated based on the continuous low-level action $a^{i}_{l}$, as defined below:
-(3.9) $\mathcal{P}^{\text{opt}}_{i,t}=\bar{a}^{i}_{l}\cdot(\mathcal{P}^{\max}_{i,t}-\mathcal{P}^{\min}_{i,t})+\mathcal{P}^{\min}_{i,t}.$
-Based on the selected action $a^{i}_{t}$ and the optimal power decision $\mathcal{P}^{\text{opt}}_{i,t}$ at each time slot, the reward function $r_{l}(s^{i}_{l},a^{i}_{l})$ (abbreviated as $r^{i}_{l}$) for the low-level agent (pile) $c_{i}$ is formulated based on the charging cost at the current time slot $t$ and the difference between the current SoC and the expected SoC at the EV’s departure connected to $c_{i}$, defined as follows:
-(3.10) $r_{l}=\omega\cdot(-\mathcal{P}^{\text{opt}}_{i,t}\cdot p_{t})+(-|SoC^{i}_{t}-SoC^{i}_{\text{dep}}|),$
+(4.9) $\mathcal{P}^{\text{opt}}_{i,t}=a^{i}_{l}\cdot(\mathcal{P}^{\max}_{i,t}-\mathcal{P}^{\min}_{i,t})+\mathcal{P}^{\min}_{i,t}.$
+Based on the optimal power decision $\mathcal{P}^{\text{opt}}_{i,t}$ at each time slot, the reward $r^{i}_{l}$ for the low-level agent (pile) $c_{i}$ is formulated considering the charging cost at the current time slot $t$ and the difference between the current SoC and the expected SoC at the EV’s departure connected to $c_{i}$, as defined below:
+(4.10) $r^{i}_{l}=\omega\cdot(-\mathcal{P}^{\text{opt}}_{i,t}\cdot p_{t})+(-|SoC^{i}_{t}-SoC^{i}_{\text{dep}}|),$
 where $p_{t}$ and $SoC^{i}_{t}$ are the energy price and the SoC of the EV at time slot $t$, respectively.
@@
 The parameter $\omega$ reflects the impact of the current charging cost.
+This reward design encourages low-level agents to discharge EVs when the energy price is high, while also aiming to meet their expected SoC.
 4.3 Optimization for the Hierarchical Control.
-Based on Sec.
-3.1 and Sec.
-3.2, the critic and actor networks are optimized with minibatches of samples from the replay buffers, which follows the soft update rule.
-Let $\theta_{h}=\{\theta^{Q}_{h},\theta^{\mu}_{h}\}$ and $\theta^{i}_{l}=\{\theta^{Q,i}_{l},\theta^{\mu,i}_{l}\}$ represent the parameters of target critic and actor networks of the high-level agent and the low-level agent ($c_{i}$), respectively.
-The parameters updates are as follows:
+In Secs.
+4.1 and 4.2, the critic and actor networks are optimized using mini-batches sampled from the replay buffers.
+To stabilize training, a soft target-network update is applied.
+Let $\theta_{h}=\{\theta^{Q}_{h},\theta^{\mu}_{h}\}$ and $\theta^{i}_{l}=\{\theta^{Q,i}_{l},\theta^{\mu,i}_{l}\}$ denote the parameters of the high-level and the $i$-th low-level networks, respectively, and let $\theta_{h}^{\prime}=\{\theta^{\prime Q}_{h},\theta^{\prime\mu}_{h}\}$ and $\theta^{{}^{\prime}i}_{l}=\{\theta^{\prime Q,i}_{l},\theta^{\prime\mu,i}_{l}\}$ be their corresponding target network parameters.
+These target parameters are updated as follows:
 (4.11) $\theta^{\prime}_{h}\leftarrow\tau\theta_{h}+(1-\tau)\theta^{\prime}_{h},\quad\theta^{{}^{\prime}i}_{l}\leftarrow\tau\theta^{i}_{l}+(1-\tau)\theta^{{}^{\prime}i}_{l},$
 where $\tau\in[0,1]$ is the coefficient controlling the update rate.
-Through these optimizations, the decision-making processes of both high-level and low-level agents are refined without requiring future information.
-The details of optimization formulation are given in Appendix A.3.
+The overall architecture of HUCA is shown in Algorithm 3.
+This design enables both high-level and low-level agents to learn optimal charging control policies by considering dynamic factors.
+Therefore, HUCA provides more effective real-time charging control without requiring future information.
+Algorithm 3 The HUCA framework
+1:The maximum learning episode $E$; charging piles $\mathcal{C}=\{c_{1},c_{2},...,c_{N}\}$; EV set $\mathcal{V}_{t}$; EV charging information $\{\mathcal{I}^{i}|v_{i}\in\mathcal{V}_{t}\}$; power boundaries $\{\mathcal{P}^{\text{max}}_{i,t},\mathcal{P}^{\text{min}}_{i,t}|v_{i}\in\mathcal{V}_{t}\}$.
+2:Optimal policies $\mu_{h}$ and $\{\mu_{l}^{\,i}\}_{i=1}^{N}$ with parameters $\theta^{\mu}_{h}$ and $\{\theta^{\mu,i}_{l}\}_{i=1}^{N}$
+3:Initialize parameters $\theta^{\mu}_{h},\theta^{Q}_{h}$ for the high-level agent and $\{\theta^{\mu,i}_{l},\theta^{Q,i}_{l}\}_{i=1}^{N}$ for the low-level agents; replay buffers $\mathcal{D}_{h},\mathcal{D}_{l}$.
+4:for episode = 1 to $E$ do
+5:  Receive initial state $s_{h}$, $\textbf{x}=\{s^{i}_{l}\}_{i=1}^{N}$
+6:  for $t=1$ to max-episode-length do
+7:    // Select actions
+8:    Select and compute action $a^{disc}_{h}$
+9:    for each low-level agent i=1,,Ni=1,...,N do
+10:     Select and compute action $a^{i}_{l}$ based on $a^{disc}_{h}$
+11:     Estimate optimal power $\mathcal{P}^{\text{opt}}_{i,t}$
+12:    end for
+13:    // Store transition experiences
+14:    Apply optimal powers $\{\mathcal{P}^{\text{opt}}_{i,t}\}_{i=1}^{N}$ to EVs and receive rewards $(r_{h},\{r^{i}_{l}\}_{i=1}^{N})$
+15:    Observe next states $(s^{\prime}_{h},\textbf{x}^{\prime}=\{s^{\prime i}_{l}\}_{i=1}^{N})$
+16:    Store $(s_{h},a_{h},r_{h},s^{\prime}_{h})$ in $\mathcal{D}_{h}$
+17:    Store $(\textbf{x},a^{1}_{l},...,a^{N}_{l},r^{1}_{l},...,r^{N}_{l},\textbf{x}^{\prime})$ in $\mathcal{D}_{l}$
+18:    Update current states
+19:  end for
+20:  // Update networks
+21:  Sample a mini-batch $\mathcal{B}_{h}$ from $\mathcal{\mathcal{}}{D}_{h}$
+22:  UpdateHigh($\mathcal{B}_{h}$, $\theta^{Q}_{h}$, $\theta^{\mu}_{h}$) $\triangleright$ Algorithm 1
+23:  Sample a mini-batch $\mathcal{B}_{l}$ from $\mathcal{D}_{l}$
+24:  for each low-level agent i=1,,Ni=1,...,N do
+25:    UpdateLow($\mathcal{B}_{l}$, $\theta^{Q,i}_{l}$, $\theta^{\mu,i}_{l}$) $\triangleright$ Algorithm 2
+26:  end for
+27:  Soft target networks update
+28:end for
 5 Experimental Results
-4.1 Dataset and Experimental Setup.
-4.1.1 Dataset Description and Preprocessing.
-For our workplace charging system scenario, we use the following datasets: (i) Building electricity demands: the CU-BEMS dataset [17] provides office building electricity consumption.
-(ii) Pricing mechanism: the ComEd APIs [5] supply real-time pricing data.
-Also, the penalty mechanism is designed based on [18].
-The data is aggregated into hourly intervals.
-The training period is from July 1, 2018, to August 31, 2018, with testing in the following month.
-Details are in Appendix A.1.
-4.1.2 Simulation of EV Behaviors.
-The charging information $\mathcal{I}^{i}=(t^{i}_{\text{arr}},t^{i}_{\text{dep}},SoC^{i}_{\text{arr}},SoC^{i}_{\text{dep}},C^{i})$ (Def.
-2) of the EV is modeled using truncated normal distributions by following [10].
-Table 1 shows the setting of random variables and $C^{i}$ is set to 60 kWh.
-Two departure scenarios are examined:
-  1. 1.
-    Certain departure scenario: The actual departure time $\widehat{t}^{i}_{dep}$ is the same as the expected time $t^{i}_{dep}$.
-  2. 2.
-    Uncertain departure scenario: The actual departure time is randomly sampled earlier than the expected time, $\widehat{t}^{i}_{dep}\in[1,t^{i}_{dep})$.
+5.1 Dataset and Experimental Setup
+5.1.1 Dataset and Preprocessing.
+For our workplace charging scenario, following datasets are used: (i) Building electricity demands: the CU-BEMS dataset [Pipattanasomporn2020] provides office building electricity consumption recorded at one-minute intervals from July 2018 to December 2019.
+(ii) Pricing mechanism: we use the ComEd APIs [ComEdHourlyPricing], which provide real-time 5-minute electricity pricing data through its hourly pricing program.
+All data are aggregated into hourly intervals.
+The training period spans July 1 to August 31, 2018, with testing conducted in the following month.
+5.1.2 Penalty Mechanism.
+The penalty charge is typically two or three times the basic capacity rate [rosado2020framework].
+We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption) [Fernndez2013CostOO] [rosado2020framework], with detailed settings shown in Table 1.
+Let the maximum instant electricity load during the testing period be $L_{max}$.
+If $L_{max}<C_{contract}$, then the penalty is set to zero.
+Given a threshold coefficient $0<\delta<1$, the penalty is calculated by dividing the excess electricity into two tiers:
+  1. (a)
+    Tier 1 (Up to 10% Overload): For overloads up to 10% of the contract capacity ($0<L_{max}\leq\delta\cdot C_{contract}$), the penalty charge is set to $2\cdot R_{base}$.
+  2. (b)
+    Tier 2 (Beyond 10% Overload): For overloads exceeding 10% of the contracted capacity ($L_{max}>\delta\cdot C_{contract}$), the penalty charge is set to $3\cdot R_{base}$ for the amount that surpasses the 10% threshold.
+Based on this setting, the total penalty cost is estimated as:
+(5.12) $\hskip 0.0pt\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
+$\hskip 71.13188pt+3\cdot R_{base}\cdot\max(0,L_{max}-\delta\cdot C_{contract}).$
+Table 1: Design of the penalty mechanism.
+Parameter Value
+Contract Capacity ($C_{contract}$) 700 kW
+Basic Capacity Rate ($R_{base}$) \$15/kW/month
+Exceed Electricity Amount Penalty Charge
+Up to 10% over $C_{contract}$ $2\cdot R_{base}$
+More than 10% over $C_{contract}$ $3\cdot R_{base}$
 Table 2: Random variables for EV information.
@@
 Expected SoC at departure $\mathcal{N}(0.8,0.1^{2})$ $0.6\leq SoC^{i}_{\text{dep}}\leq 0.9$
-Table 2: Performance comparison (HUCA performance) with different number of charging piles (CP).
-Except of the “OPT” method, the best and second best results for “Penalty Cost” ($\downarrow$) and “Total Cost” ($\downarrow$) are in bold and underlined, respectively.
-Cells in gray highlight the results that are greater than or equal to the median value for “SoC Maintenance” ($\uparrow$), “SoC Fulfillment” ($\uparrow$) and “User Satisfaction” ($\uparrow$).
-Certain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
-Penalty Cost (USD) 10 0.00 3682.85 2516.84 522.28 354.84 354.82
-Total Cost (USD) 5805.28 9563.31 8426.36 6431.31 6220.33 6215.42
-Penalty Cost (USD) 20 0.00 7818.52 5643.15 2271.32 1869.38 1643.47
-Total Cost (USD) 5860.05 13929.11 11798.68 8413.61 7952.69 7728.57
-Uncertain Departure Scenario Metric CP OPT DDPG IQL VDN MADDPG HUCA
-Penalty Cost (USD) 10 0.00 4671.74 1742.35 529.18 448.52 355.84
-Total Cost (USD) 5691.96 10398.77 7479.04 6214.53 6087.48 5993.78
-SoC Fulfillment (%) 53.77 54.05 46.04 40.34 27.10 40.77
-SoC Maintenance (%) 42.52 -10.58 129.01 -120.61 11.87 45.40
-User Satisfaction (%) 48.14 21.73 87.52 -40.13 19.48 43.08
-Penalty Cost (USD) 20 0.00 5010.81 2002.12 966.51 931.01 354.90
-Total Cost (USD) 5760.05 10823.07 7815.99 6721.82 6562.80 6013.52
-SoC Fulfillment (%) 56.06 55.27 47.94 48.88 40.34 42.44
-SoC Maintenance (%) 32.77 12.56 0.90 91.96 -1.16 124.41
-User Satisfaction (%) 44.41 33.91 24.42 70.42 19.59 83.42
-Table 3: Ablation study results for the uncertain departure scenario. (SoC fulfillment and maintenance scores are abbreviated as “Ful.” and “Main.”, respectively.)
-CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
-10 HUCA 355.84 5993.78 40.77 % 45.40%
-w/o C.A. 2015.62 7761.55 62.49% 124.01%
-w/o H. 2707.71 8483.96 63.10% 31.91%
-w/o Either 723.46 6393.26 44.53% -12.53%
-CP Method Penalty $\downarrow$ Total $\downarrow$ Ful. $\uparrow$ Main. $\uparrow$
-20 HUCA 354.90 6013.52 42.44% 124.41%
-w/o C.A. 9362.87 15309.23 69.73% 209.70%
-w/o H. 3461.37 9261.72 58.61% 7.69%
-w/o Either 618.46 6381.87 53.26% 43.02%
+[…]
+Certain Departure Scenario
+Metrics CP Num. Methods
+OPT DDPG IQL VDN MADDPG HUCA (ours)
+Penalty Cost ($\downarrow$) 10 0.00(▲100.0%) 3682.85(▽937.9%) 2516.84(▽609.3%) 522.28(▽47.2%) 354.84(▽0.0%) 354.82
+Total Cost ($\downarrow$) 5805.28(▲6.6%) 9563.31(▽53.9%) 8426.36(▽35.6%) 6431.31(▽3.5%) 6220.33(▽0.1%) 6215.42
+Penalty Cost ($\downarrow$) 20 0.00(▲100.0%) 7818.52(▽375.7%) 5643.15(▽243.4%) 2271.32(▽38.2%) 1869.38(▽13.7%) 1643.47
+Total Cost ($\downarrow$) 5860.05(▲24.2%) 13929.11(▽80.2%) 11798.68(▽52.7%) 8413.61(▽8.9%) 7952.69(▽2.9%) 7728.57
+Uncertain Departure Scenario
+Metrics CP Num. Methods
+OPT DDPG IQL VDN MADDPG HUCA (ours)
+Penalty Cost ($\downarrow$) 10 0.00(▲100.0%) 4671.74(▽1212.9%) 1742.35(▽389.6%) 529.18(▽48.7%) 448.52(▽26.0%) 355.84
+Total Cost ($\downarrow$) 5691.96(▲5.0%) 10398.77(▽73.5%) 7479.04(▽24.8%) 6214.53(▽3.7%) 6087.48(▽1.6%) 5993.78
+[…]
+[…]
+[…]
+Penalty Cost ($\downarrow$) 20 0.00(▲100.0%) 5010.81(▽1311.6%) 2002.12(▽464.1%) 966.51(▽172.3%) 931.01(▽162.3%) 354.90
+Total Cost ($\downarrow$) 5760.05(▲4.2%) 10823.07(▽80.0%) 7815.99(▽30.0%) 6721.82(▽11.8%) 6562.80(▽9.1%) 6013.52
+[…]
+[…]
+[…]
+5.1.3 Simulation of EV Behaviors.
+The charging information $\mathcal{I}^{i}=(t^{i}_{\text{arr}},t^{i}_{\text{dep}},SoC^{i}_{\text{arr}},SoC^{i}_{\text{dep}},C^{i})$ (Def.
+2) of EV $v_{i}$ is modeled using normal distributions by following [LiConstrainedLR].
+Table 2 shows the setting of random variables and $C^{i}$ is set to 60 kWh.
+Two departure scenarios for determining the actual departure time $\widehat{t}^{i}_{dep}$ of EV $v_{i}$ are examined:
+  1. (a)
+    Certain departure scenario: $\widehat{t}^{i}_{dep}$ is the same as the expected departure time $t^{i}_{dep}$.
+  2. (b)
+    Uncertain departure scenario: $\widehat{t}^{i}_{dep}$ is randomly sampled earlier than the expected departure time, $\widehat{t}^{i}_{dep}\in[1,t^{i}_{dep})$.
 5.1.4 Baselines.
-We compare our HUCA with following baselines.
-OPT: An oracle with full knowledge of future information serves the optimal solution of minimal total cost and minimal penalty cost.
-We include the single-agent (DDPG [13]) and multi-agent reinforcement learning models (IQL [22], VDN[21], and MADDPG [14]).
-These details are provided in Appendix A.1.4.
+We compare our HUCA against several baselines, including an oracle with full knowledge of future information, the single-agent model(DDPG [lillicrap2015continuous]) and multi-agent reinforcement learning models (IQL [Tan1997MultiAgentRL], VDN[Sunehag2017ValueDecompositionNF], and MADDPG [lowe2017multi]), as described below:
+  • OPT: Assumes complete knowledge of future information, including all EV charging schedules, building energy demands, and dynamic electricity prices.
+This baseline uses a linear programming optimization model to solve the charging control problem.
+  • DDPG (Deep Deterministic Policy Gradient) [lillicrap2015continuous]: An actor-critic algorithm designed for continuous control in single-agent environments.
+  • IQL (Independent Q-Learning) [Tan1997MultiAgentRL]: A decentralized multi-agent reinforcement learning approach where each agent independently learns its own Q-function.
+  • VDN (Value Decomposition Networks) [Sunehag2017ValueDecompositionNF]: A centralized value-based reinforcement learning method used in multi-agent systems.
+  • MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [lowe2017multi]: Extends DDPG to multi-agent settings by incorporating centralized training and decentralized execution.
+Note that for all models (except for OPT), the states, actions, and rewards are configured in the same manner as for our low-level agents, but without incorporating the information provided by the high-level agent network.
 5.1.5 Evaluation Metrics.
-  • Penalty Cost (USD) measures the cost of exceeding the contracted capacity [7], which multiplies the penalty rate with the peak amount of electricity (check Appendix A.1 for details).
+Following metrics are used to evaluate performance:
+  • Penalty Cost (USD) measures the cost of exceeding the contracted capacity [Fernndez2013CostOO], which multiplies the penalty rate with the peak amount of electricity.
+In accordance with [rosado2020framework], the penalty mechanism is designed as detailed in Table 1.
   • Total Cost (USD) is the sum of the basic electricity cost (electricity load $L_{t}$ multiplied by dynamic pricing $p_{t}$) and the penalty cost over the testing data.
 In the certain departure scenario, all EV charging demands are met since all baselines are subject to the charging power limitations (Sec.
-2.1, Def.
+3.1, Def.
 3) and no unexpected EV departures occur.
@@
 The performance of these three metrics is reported as the average for all EVs per parking session.
+5.1.6 Hyperparamter Settings.
+The charging capacity ($\mathcal{P}^{\max}$) is set to 150 (kW), the EV battery capacity $C^{i}$ (Def.
+2) is 60 (kwh), and the charging efficiency is 95%.
+For training our HUCA model, we use 1,500 episodes, with a replay buffer capacity of 30,000 for each low-level agent, and a batch size of 1,024.
+The penalty coefficient $\varphi$ (Eq.
+4.2) weight factors $\kappa$ (Eq.
+4.2), $\omega$ (Eq.
+4.10), and the importance of the uncertainty term $\rho$ (Eq.
+4.6) are set to 0.1, 0.1, 0.5, and 10, respectively.
+5.2 Comparison Results.
+As shown in Table 3, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total cost without relying on future information (such as future energy prices, EV charging requests, and actual EV departure times) for charging control across both departure scenarios.
+In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs.
+While the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, its performance exceeds the median of the baselines in most cases (highlighted in gray).
+In contrast, although DDPG and IQL achieve higher SoC fulfillment under different numbers of charging piles, they incur substantially higher total costs than HUCA (by 20-80%).
+Combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
+Among multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost.
+However, user satisfaction scores of MADDPG are much lower than those of HUCA.
+While VDN achieves higher SoC fulfillment than HUCA with 20 charging piles, its maintenance scores are significantly lower.
+In some cases, the maintenance scores of MADDPG and VDN fall into negative values, indicating that EVs primarily serve as power providers and return with lower SoCs than at arrival, leading to additional expenses for users.
+[…]
+Uncertain Departure Scenario
+CP Method Penalty ($\downarrow$) Ful. ($\uparrow$) Main. ($\uparrow$)
+10 HUCA 355.84 40.77 45.40
+w/o C.A. 2015.62(▽466.4%) 62.49(▲53.3%) 124.01(▲173.1%)
+w/o H. 2707.71(▽660.9%) 63.10(▲54.8%) 31.91(▽29.7%)
+w/o Either 723.46(▽103.3%) 44.53(▲9.2%) -12.53(▽127.6%)
+CP Method Penalty ($\downarrow$) Ful. ($\uparrow$) Main. ($\uparrow$)
+20 HUCA 354.90 42.44 124.41
+w/o C.A. 9362.87(▽2538.2%) 69.73(▲64.3%) 209.70(▲68.6%)
+w/o H. 3461.37(▽875.2%) 58.61(▲38.1%) 7.69(▽93.8%)
+w/o Either 618.46(▽74.3%) 53.26(▲25.5%) 43.02(▽65.4%)
+5.3 Ablation Study and Uncertainty Impact.
+We conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o Either).
+The results are shown in Table 4 and Figure 4.
+In Table 4, removing either the critic augmentation or the high-level agent increases SoC fulfillment and maintenance scores (improvements of about 9–173%).
+However, this also results in a much larger rise in penalty costs (worsening by about 74–2538%).
+The increase in penalty costs far outweighs the gains in SoC fulfillment and maintenance when removing each component.
+On the other hand, removing both components results in a slight increase in costs but causes a significant drop in SoC maintenance scores.
+Figure 4 also shows that with more charging piles (CP=20), removing critic augmentation increases user satisfaction but also raises the total cost substantially (worsening by 154%) compared to the full HUCA model.
+Figure 4: Ablation study on total cost and user satisfaction (CP=20).
+These findings indicate that the hierarchical structure and critic augmentation in HUCA complement each other, and their combined use is essential for achieving the lowest costs while balancing energy supply between the building and EVs.
 Figure 5: Performance of HUCA with different $\rho$ values with 10 piles in the uncertain departure scenario.
-4.2 Comparison Results.
-As shown in Table 2, except for the ”OPT” method (best-case result), HUCA achieves the lowest penalty and total costs that do not rely on future information for charging control across both departure scenarios.
-In addition, its total cost is comparable to OPT, manifesting its effectiveness in minimizing costs.
-Although the SoC fulfillment, maintenance, and user satisfaction scores of HUCA are not the best, it still achieves a comparable results to all the other baselines.
-Note that combining the above observations, HUCA actually strikes the best balance between the trade-off of electricity costs and fulfilling user demands by introducing the hierarchical control and the uncertainty-aware critic augmentation.
-Among the multi-agent models, VDN and MADDPG come closest to HUCA in terms of cost.
-However, the user satisfaction scores of MADDPG are much lower than those of HUCA across different numbers of charging piles.
-While VDN achieves higher SoC fulfillment score than HUCA with 20 charging piles, its maintenance scores are significantly lower.
-With 10 charging piles, VDN’s maintenance score even drops to a negative value (-120.61%), indicating that under MADDPG and VDN control, EVs’ SoCs are not maintained at arrival levels.
-Therefore, EVs primarily serve as power providers rather than being effectively charged, resulting in additional expenses for users.
-4.3 Ablation Study and Analysis of the Uncertainty Impact.
-We conduct an ablation study by removing the uncertainty-aware critic augmentation (w/o C.A.), the high-level agent (w/o H.), and both of them (w/o Either).
-The results are shown in Table 3.
-Compared to HUCA, removing either the critic augmentation or the high-level agent increases SoC fulfillment and maintenance scores.
-However, this also leads to a significant rise in both penalty and total costs.
-On the other hand, removing both components results in a slight increase in costs but causes a significant drop in SoC maintenance scores.
-These findings indicate that the hierarchical structure and critic augmentation in HUCA complement each other, and their combined use is essential for achieving the lowest costs while effectively balancing energy supply between the office building and EVs.
 Furthermore, to assess the impact of the uncertainty term in HUCA for handling uncertain departures, Figure 5 compares total cost and the SoC fulfillment under various uncertainty coefficient $\rho$ (Eq.
-3.6).
+4.6).
 As $\rho$ increases, total cost decreases significantly.
@@
 Properly tuning $\rho$ can further improve this balance, enhancing the overall effectiveness of HUCA.
-4.4 Case Study.
-To examine how HUCA balances energy supply between the building and EVs, Figure 4 visualizes the control decisions over a day.
-Two significant EV charging behaviors appear at 7:00 and 16:00 (red rectangle).
-At 7:00, both electricity demand and prices for the building are minimal (Figure 4(a)), therefore HUCA focuses on charging the EVs to reach their target SoC.At 16:00, even though there is a high demand for building energy and elevated prices, which typically would lead to transferring energy from the EVs to the building to reduce costs, HUCA opts to charge the EVs because of their high charging urgency (Figure 4(b)) using Eq.
-(3.6).
-In contrast, at 8:00 (green rectangle), HUCA discharges EVs to serve the high-usage building and avoid high pricing time simultaneously.
-These results manifest that HUCA effectively and dynamically balances energy supply between the building and EVs based on real-time information.
 Figure 6: Details of the charging control for a day from 7:00 to 17:00.
 (a) shows building electricity demands and dynamic pricing; (b) visualizes average charging urgency, with $\Delta SoC$ as the difference between current and expected SoC, and $\Delta T$ as the remaining time until planned departure; (c) presents final control decisions.
-5 Related Work
-5.1 Real-time Charging Control.
-Guo et al.
-[8] combined day-ahead optimization with real-time control to address uncertainties in solar output and EV parking behaviors.
-Chandra Mouli et al.
-[16] introduced an energy management system with vehicle-to-grid (V2G) services to adjust charging control.
-Chopra et al.
-[3] conducted a cost-benefit analysis for EV charging infrastructure in parking lots.
-Zhao et al.
-[29] employed risk-based day-ahead scheduling for EV aggregators using information gap decision theory to minimize building costs.
-Zhang et al.
-[28] developed an EV load optimization model to reduce grid-level energy costs by considering peak and valley electricity prices.
-Chen et al.
-[2] devised strategies for battery energy storage systems to participate in peak load shifting, minimizing load fluctuations.
-5.2 Reinforcement Learning for Charging Management.
-Single-agent RL, multi-agent RL (MARL), and hierarchical MARL approaches were used to manage charging systems.
-Single-agent RL represented the control system [12, 15] under various considerations, such as ensuring EV battery requirements [26] and addressing charger shortage [9].
-MARL modeled fine-grained cooperations [25] among EVs and buildings to tackle safety concerns [27] and power transfer overload [6].
-Hierarchical MARL further separated system operators and EV users, and coordinated them to minimize demand charges and energy costs [20].
-Although literature in Sec.
-5.1 and  5.2 optimized charging control with various aspects, most studies do not address the issue of uncertain EV departures.
-While some works[15, 8, 27] considered uncertainties in EV behavior, they primarily focused on a single objective, such as maximizing system revenue or minimizing user charging costs.
-In contrast, our work tackles dynamic factors, including fluctuating electricity prices and uncertain EV departures, aiming to simultaneously minimize total electricity costs for the charging system while accounting for EV users’ charging demands.
-6 Conclusion
-In this paper, we propose HUCA, a novel hierarchical multi-agent framework designed for real-time charging control in the workplace charging scenario.
-To address practical dynamics and limitations, HUCA integrates a hierarchical control and uncertainty-aware critic augmentation.
-With this design, HUCA adapts to dynamic factors, optimizes charging or discharging power for each charging pile, and accounts for departure uncertainties in evaluating the quality of charging control.
-Experiments demonstrate that, in both certain and uncertain departure scenarios, HUCA outperforms existing methods by achieving the lowest electricity costs while maintaining competitive performance in meeting EV charging requirements.
+5.4 Case Study.
+To examine how HUCA balances energy supply between the building and EVs, Figure 6 visualizes the control decisions over a single day for all EVs, which can be discussed in terms of charging and discharging behaviors:
+Charging Behaviors.
+Two significant EV charging behaviors appear at 7:00 and 16:00 (red rectangle) but with different reasons.
+At 7:00, both electricity demand and prices for the building are minimal (Figure 6(a)), therefore HUCA focuses on charging the EVs to reach their target SoC.
+At 16:00, even though there is a high demand for building energy and elevated prices, which typically would lead to transferring energy from the EVs to the building to reduce costs, HUCA opts to charge the EVs because of their high charging urgency (Figure 6(b)).
+Discharging Behaviors.
+A similar situation occurs at 8:00 (green rectangle); however, since the charging urgency is relatively low at this time, HUCA discharges the EVs to serve the high-usage building while simultaneously avoiding high-priced electricity usage.
+These results manifest that HUCA effectively and dynamically balances energy supply between the building and EVs based on real-time information.
+6 Conclusion and Future Work
+In this paper, we propose HUCA, a novel hierarchical multi-agent framework designed for real-time charging control.
+To address practical dynamics and limitations, HUCA integrates hierarchical control and uncertainty-aware critic augmentation to adapt to dynamic factors, optimize charging decisions, and account for departure uncertainties.
+Experiments show that HUCA outperforms existing methods in both certain and uncertain departure scenarios.
 A case study further illustrates that HUCA effectively balances energy supply between the building and EVs using real-time information.
-A Reproducibility
-A.1 Detailed Experimental Settings.
-A.1.1 Dataset Preprocessing.
-We utilize the CU-BEMS dataset [17], which contains detailed energy consumption data from an office building, recorded at one-minute intervals from July 2018 to December 2019.
-For dynamic pricing (real-time pricing) data, we use the APIs [5] provided by ComEd (Commonwealth Edison), an electric utility company, offering real-time 5-minute electricity pricing data through its hourly pricing program.
-In our study, we aggregate the total energy consumption from all electrical appliances in the CU-BEMS dataset into hourly data to represent the building’s energy consumption in our workplace charging scenario.
-The dynamic pricing data obtained from ComEd APIs is similarly aggregated into hourly intervals.
-The training period covers July 1, 2018, to August 31, 2018, and the testing period runs from September 1, 2018, to September 30, 2018.
-A.1.2 Penalty Mechanism.
-For the design of the penalty mechanism, the penalty charge is typically two or three times the basic capacity rate [18].
-We design the penalty system based on the amount of electricity exceeding the contracted capacity (the upper bound of instant electricity consumption)[7, 18], with detailed settings shown in Table 4.
-Let the maximum instant electricity load during the testing period be $L_{max}$.
-If $L_{max}<C_{contract}$, then the penalty is set to zero.
-Given a threshold coefficient $10<\delta<$ 1, the penalty is calculated by dividing the excess electricity into two tiers:
-  1. 1.
-    Tier 1 (Up to 10% Overload): For overloads up to 10% of the contract capacity ($0<L_{max}\leq\delta\cdot C_{contract}$), the penalty charge is set to $2\cdot R_{base}$.
-  2. 2.
-    Tier 2 (Beyond 10% Overload): For overloads exceeding 10% of the contracted capacity ($L_{max}>\delta\cdot C_{contract}$), the penalty charge is set to $3\cdot R_{base}$ for the amount that surpasses the 10% threshold.
-Table 4: Design of the penalty mechanism.
-Parameter Value
-Contract Capacity ($C_{contract}$) 700 kW
-Basic Capacity Rate ($R_{base}$) \$15/kW/month
-Exceed Electricity Amount Penalty Charge
-Up to 10% over $C_{contract}$ $2\cdot R_{base}$
-More than 10% over $C_{contract}$ $3\cdot R_{base}$
-Based on this setting, the total penalty cost is estimated as follows:
-(A.1) Penalty $\text{Penalty Cost}=2\cdot R_{base}\cdot\min(L_{max},\delta\cdot C_{contract})$
-$\hskip 71.13188pt+3\cdot R_{base}\cdot\max(0,L_{max}-\delta\cdot C_{contract}).$
-A.1.3 Hyperparamter Settings.
-The charging capacity ($\mathcal{P}^{\max}$) is set to 150 (kW), the EV battery capacity $C^{i}$ (Def.
-2) is 60 (kwh), and the charging efficiency is 95%.
-For training our HUCA model, we use 1,500 episodes, with a replay buffer capacity of 30,000 for each low-level agent, and a batch size of 1,024.
-The penalty coefficient $\varphi$ (Eq.
-3.2) weight factors $\kappa$ (Eq.
-3.2), $\omega$ (Eq.
-3.10), and the importance of the uncertainty term $\rho$ (Eq.
-3.6) are set to 0.1, 0.1, 0.5, and 10, respectively.
-A.1.4 Baseline Settings.
-To evaluate the performance of the proposed HUCA model, we compared it against several established baseline approaches:
-  • OPT: Assumes complete knowledge of future information, including all EV charging schedules, building energy demands, and dynamic electricity prices.
-This baseline uses a linear programming optimization model to solve the charging control problem.
-  • DDPG (Deep Deterministic Policy Gradient) [13]: An actor-critic algorithm designed for continuous control in single-agent environments.
-  • IQL (Independent Q-Learning) [22]: A decentralized multi-agent reinforcement learning approach where each agent independently learns its own Q-function.
-  • VDN (Value Decomposition Networks)[21]: A centralized value-based reinforcement learning method used in multi-agent systems.
-  • MADDPG (Multi-Agent Deep Deterministic Policy Gradient) [14]: Extends DDPG to multi-agent settings by incorporating centralized training and decentralized execution.
-Note that for all models (except for OPT), the states, actions, and rewards are configured in the same manner as for our low-level agents, but without incorporating the information provided by the high-level agent network.
-A.2 State representations of high-level and low-level MDPs.
-A.2.1 High-level Agent.
-A state $s_{h}$ in the MDP optimizied by the high-level agent includes three main information $s_{h}=(I^{e}_{t},I^{v}_{t},I^{l}_{t-1})$ at current time slot $t$. The details are as follows:
-  • $I^{e}_{t}$ includes the electricity load of the building and electricity price for the following periods: the current time slot, the average over the past $n$ hours, and the historical price at the same time and day of the week.
-  • $I^{v}_{t}$ includes the current number of EVs at the charging station and the amount of energy provided to them.
-  • $I^{l}_{t-1}$ includes the average critic value (from the action-value function) and the standard deviation of all low-level agents from the previous time slot.
-A.2.2 Low-level Agent.
-For a low-level agent $c_{i}\in\mathcal{C}$, the state $s^{i}_{l}$ observed at time slot $t$ can be described as comprising two main information $s^{i}_{l}=(I^{v_{j}}_{t},I^{h}_{t})$. Details are:
-  • $I^{v_{j}}_{t}$ includes the current SoC of EV $v_{j}$ and the suitable maximum and minimum charging powers ($\mathcal{P}^{\max}_{i,t}$ and $\mathcal{P}^{\min}_{i,t}$) of agent (charging pile) $c_{i}$ as defined in Eq.
-2.1.
-  • $I^{h}_{t}$ includes the discrete high-level action $a^{\text{disc}}_{h}$ determined at time slot $t$, and the critic value of state-action pair $(s_{h},a_{h})$ obtained from the high-level critic network $Q_{h}(s_{h},a_{h};\theta^{Q}_{h})$ at time slot $t$.
-A.3 Optimization of HUCA.
-We optimize the HUCA framwork using minibatches of samples from the replay buffers.
-For the high-level agent, given a minibatch of $\mathcal{M}$ samples from the replay buffer $\mathcal{D}_{h}$, the training loss of the critic network is rewritten as:
-(A.2) $\mathcal{L}(\theta^{Q}_{h})=\frac{1}{\mathcal{M}}\sum_{m}\Bigl{(}Q_{h}(s^{m}_{h},a^{m}_{h})-y^{m}_{h}\Bigr{)}^{2}.$
-The actor network is updated using the sampled policy gradient as:
-(A.3) $\nabla_{\theta^{\mu}_{h}}J(\mu_{h})\approx\frac{1}{\mathcal{M}}\sum_{m}\nabla_{\theta^{\mu}_{h}}\mu_{h}(a^{m}_{h}|s^{m}_{h})\nabla_{a_{h}}Q_{h}(s^{m}_{h},a^{m}_{h}).$
-Similarly, the critic and actor networks of each low-level agent $c_{i}$ are also optimized using minibatches.
-B Limitation and Future Work
-Despite the strong performance of HUCA and its consideration of multiple dynamic factors, we have not yet accounted for user willingness to adjust their charging preferences to supply energy to the building.
-A promising direction for future work is to incorporate user willingness, along with the buy-and-sell energy behaviors between users and the system, into our operational framework.
-User willingness could be integrated into the low-level actor-critic networks of HUCA.
-In our current design, the departure uncertainty term is used to set a lower confidence bound for the action-value function of each low-level agent.
-When incorporating user willingness, an incentive mechanism could be developed to estimate the upper confidence bound of the action-value function, as user willingness and incentives might increase the effectiveness of charging control.
-This would not only help identify which EV users have greater flexibility in participating in energy control but also allow the system to evaluate both the lower and upper bounds of the potential quality of low-level actions.
-This approach could improve the optimal charging and discharging power control, enhancing both user satisfaction and the overall effectiveness of the system.
+One line of the future work is to account for user willingness when adjusting charging preferences to supply energy to the building, e.g., with both user willingness and buy-and-sell energy behaviors in HUCA.
+Incorporating them can help improve the estimation of the confidence bound for the action-value function, leading to a more efficient charging control.